### v1

# Solving Inverse Problems with FLAIR

Abstract
Flow-based latent generative models such as Stable Diffusion 3 are able to generate images with remarkable quality, even enabling photorealistic text-to-image generation. Their impressive performance suggests that these models should also constitute powerful priors for inverse imaging problems, but that approach has not yet led to comparable fidelity. There are several key obstacles: (i) the encoding into a lower-dimensional latent space makes the underlying (forward) mapping non-linear; (ii) the data likelihood term is usually intractable; and (iii) learned generative models struggle to recover rare, atypical data modes during inference. We present FLAIR, a novel training free variational framework that leverages flow-based generative models as a prior for inverse problems. To that end, we introduce a variational objective for flow matching that is agnostic to the type of degradation, and combine it with deterministic trajectory adjustments to recover atypical modes. To enforce exact consistency with the observed data, we decouple the optimization of the data fidelity and regularization terms. Moreover, we introduce a time-dependent calibration scheme in which the strength of the regularization is modulated according to off-line accuracy estimates. Results on standard imaging benchmarks demonstrate that FLAIR consistently outperforms existing diffusion- and flow-based methods in terms of reconstruction quality and sample diversity.
1 Introduction
Flow-based generative models are at the core of modern image generators like Stable Diffusion or FLUX esser24icml . Beyond image generation based on text prompts, these models have emerged as powerful data-driven priors for a whole range of visual computing tasks. Their comprehensive representation of the visual world, learned from internet-scale training datasets, makes them an attractive alternative to traditional handcrafted image priors. Often, they can be used without any task-specific retraining.
While it is evident that a model capable of generating photorealistic images should be suitable as prior (a.k.a. regularizer) for inverse imaging problems, a practical implementation faces several challenges. On the one hand, flow-based models normally operate in the lower-dimensional latent space of a variational autoencoder (VAE), which means that the forward operator (the relationship between the observed, degraded image and the desired, clean target image) is no longer linear. On the other hand, the iterative nature of the generative process means that intermediate stages are corrupted with (time-dependent) random noise. Hence, one cannot explicitly evaluate their data likelihood, which renders the data term intractable. Moreover, learned generative models tend to overly favor regions of the training distribution that have a high sample density. For test samples that fall in low-density regions, the prior will have a too strong tendency to pull towards outputs with higher a-priori likelihood, compromising fidelity to the input observations.
Here, we propose flow-based latent adaptive inference with deterministic re-noising (FLAIR), a novel training free variational framework explicitly tailored to integrate flow-based latent diffusion models into inverse problem-solving. To the best of our knowledge, FLAIR is the first scheme that combines latent generative modeling, flow matching and variational inference into a unified formulation for inverse problems. Our main contributions are
  • A novel variational objective for inverse problems with flow-matching priors.
  • Deterministic trajectory adjustments to robustly handle rare or atypical data scenarios.
  • Decoupled optimization of data and regularization terms, enabling hard data consistency.
  • A novel, time-dependent weighting scheme, calibrated via offline accuracy estimates, that adapts the regularization along the flow trajectory to match the changing reliability of the model’s predictions, ensuring robust inference.
Figure 1: Starting from the adjoint based initialization, we alternate between (i) regularizer updates via a flow-matching loss that aligns the posterior velocity $u_{t}$ with the learned velocity field $v_{\theta}$, and (ii) hard data consistency steps that project the current estimate onto the measurement manifold.
2 Related Work
Deep learning based priors. Deep learning–based methods typically follow one of two main approaches: they either directly learn an inverse mapping zb18 ; ja08 ; le17 ; becker2023neural , or aim to learn a suitable prior, either through non-generative approaches like unrolled optimization networks ham18 ; koef21 ; aga18 ; narnhofer2021bayesian or through generative models such as generative adversarial networks na19 ; bor17 ; shah2018solving , or diffusion ho2020denoising_ddpm ; song2020denoising_ddim ; song2021scorebased respectively flow based models lipman2023flow ; liu2022flow . The latter have demonstrated impressive performance in image generation tasks, sparking growing interest in leveraging them as priors for solving inverse problems, particularly through posterior sampling techniques.
Posterior sampling. Although incorporating the prior learned from a diffusion or flow-based model seems straightforward, problems arise due to the inherent time-dependent structure of diffusion models, which makes the likelihood term intractable chung2023diffusion . A variety of approaches have been proposed for diffusion-based posterior sampling, including, enforcing the trajectory to stay on the respective noise manifold chung2023diffusion ; chung2022improving , applying an SVD to run diffusion in spectral domain kawar2021snips , utilizing range-null space decomposition during the reverse diffusion process kawar2021snips , guidance by the pseudo-inverse of the forward operator song2023pseudoinverse .
Many prior methods perform well in pixel space but are difficult to apply in latent diffusion models due to VAE non-linearity or memory constraints. In order to circumvent this issue, the authors of ReSample song2024solving rely on enforcing hard data consistency through optimization and resampling during the reverse diffusion process. PSLD rout2023solving , introduces additional objectives terms in order to ensure that all gradient updates point to the same optima in the latent space. FlowChef patel2024flowchef incorporates guidance into the flow trajectory during inference, whereas FlowDPS kim2025flowdps separates the update step into two components: one for estimating the clean image and another for estimating the noise.
In contrast FLAIR follows another class of posterior sampling-based methods, which integrate diffusion priors with inverse problems by directly optimizing a variational objective that approximates the data posterior mardani2024a . This framework was recently extended by RSD zilberstein2025repulsive , which incorporates a repulsion mechanism to promote sample diversity and applied to latent diffusion models. A known issue with this type of optimization is mode collapse pooledreamfusion , which leads to blurry results for these methods. Our method targets this problem by introducing a deterministic trajectory adjustment.
3 Background
3.1 Inverse problems
In many imaging tasks, such as inpainting bertalmio2000image , super-resolution park2003super or tomographic reconstruction sidky2008image , one aims to recover a target signal $x\in\mathbb{R}^{n}$ from a distorted observation $y\in\mathbb{R}^{m}$. The observation is regarded as the result of applying a forward operator $A:\mathbb{R}^{n}\mapsto\mathbb{R}^{m}$ to the target signal, corrupted by additive Gaussian noise $\nu\in\mathbb{R}^{m}$ with standard deviation $\sigma_{\nu}$.
$y=\mathcal{A}x+\nu.$ (1)
In most practical applications, the forward operator $\mathcal{A}$ is either non-invertible or severely ill-conditioned, making (1) generally ill-posed. Variational methods solve ill-posed inverse problems by minimizing an energy functional
$\mathcal{E}(x,z)=\mathcal{D}(x,y)+\mathcal{R}(x).$ (2)
to recover the solution. Interpreted probabilistically via Bayes’ theorem, the posterior distribution $p(x|z)$ is proportional to the product $p(z|x)p(x)$. In the negative log-domain, this yields the data term $\mathcal{D}(x,z)=-\log p(z|x)$ and the regularizer $\mathcal{R}(x)=-\log p(x)$. Handcrafted priors based on regularity assumptions like sparsity rof92 ; sa01 ; chpo11 ; da04 ; ho20_ip_review have long been the standard, but have largely been replaced by deep learning-based methods in modern data-driven schemes.
3.2 Flow based priors
Models based on flow matching lipman2023flow learn a time-dependent vector field $v_{\theta}(x_{t},t)$ that continuously transforms samples from a simple initial distribution $p_{1}(x)$ to a complex target data distribution $p_{0}(x)$. Formally, this transformation is described by solving the ordinary differential equation (ODE):
$\frac{\operatorname{d}\!{}}{\operatorname{d}\!{t}}\psi_{t}(x)=v_{\theta,t}(\psi_{t}(x)),\quad t\in[0,1],$ (3)
where $\psi_{t}(x)$ represents the trajectory of a sample evolving smoothly from an initial sample drawn at $t=1$ toward a target sample at $t=0$. Since the integrated ODE path maps the simple distribution $p_{1}(x)$ to the complex target $p_{0}(x)$, the learned flow-based model captures the structure of the data and can therefore serve as a powerful prior for solving inverse problems. To make this approach tractable for high-resolution data, we adopt the latent diffusion model (LDM) framework rombach2022high , which shifts the generative process to a lower-dimensional latent space using a pretrained autoencoder with encoder $E:\mathbb{R}^{n}\mapsto\mathbb{R}^{d}$ and decoder $D:\mathbb{R}^{d}\mapsto\mathbb{R}^{n}$ with $d\ll n$. However, applying such priors to inverse problems introduces challenges, as the non-linearity of the VAE disrupts the linear relationship between measurements and the target signal, resulting in a nonlinear forward operator.
3.3 Variational flow sampling
To solve inverse problems from a Bayesian perspective, we aim to sample from the posterior
$p(x_{0}|y)\propto p(y|x_{0})p(x_{0}),$ (4)
where the likelihood is given by $p(y|x_{0})=\mathcal{N}(\mathcal{A}x_{0},\sigma^{2}\mathrm{Id})$, and $p(x_{0})$ represents the prior modeled by the flow-based generative model. Inspired by previous work mardani2024a ; zilberstein2025repulsive we introduce a variational distribution $q(x_{0}|y)=\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ to approximate the true posterior $p(x_{0}|y)$, by minimizing their Kullback–Leibler divergence:
$q(x_{0}|y)\in\arg\min_{q(x_{0}|y)}\text{KL}(q(x_{0}|y)\|p(x_{0}|y)).$ (5)
Rewriting the KL divergence by means of the variational lower bound leads to:
$KL(q(x_{0}|y)\|p(x_{0}|y))=-\underbrace{\mathbb{E}_{q(x_{0}|y)}[\log p(y|x_{0})]}_{\text{$\mathcal{D}(x,y)$}}+\underbrace{KL(q(x_{0}|y)\|p(x_{0}))}_{\text{$\mathcal{R}(x)$}}+\underbrace{\log p(y)}_{\text{const}}.$ (6)
Since a single Gaussian cannot capture a multi-modal posterior, we simplify to a deterministic approximation, setting $\sigma_{x}^{2}=0$. Equivalently, this corresponds to a single-particle approximation in the sense of Stein variational methods liu2016stein . As shown in song2021maximum , rewriting Equation 6 under this approximation and extending it to the time-dependent noisy posterior yields:
$\arg\min_{q(x_{0}|y)}\underbrace{\mathbb{E}_{q(x_{0}|y)}\left[\frac{\|y-f(\mu_{x})\|^{2}}{2\sigma_{\nu}^{2}}\right]}_{\text{$\mathcal{D}(x,y)$}}+\underbrace{\int_{0}^{T}\omega(t)~{}\mathbb{E}_{q(x_{t}|y)}\left[\left\|\nabla_{x}\log q(x_{t}|y)-\nabla_{x}\log p(x_{t})\right\|^{2}\right]dt}_{\text{$\mathcal{R}(x)$}}$ (7)
The first term in Equation 7 describes the data term $\mathcal{D}(x,y)$ and the second the regularizer $\mathcal{R}(x)$, where the integral ensures optimization over the entire diffusion trajectory. Notably, the latter constitutes a weighted score-matching objective, where $\nabla_{x}\log p(x_{t})$ represents the score function song2021scorebased , which may be extracted from a pretrained diffusion or flow model.
The score of the noisy variational distribution depends on the forward diffusion process and can be computed analytically. Note that for $\omega(t)=\beta(t)/2$ the weighted score-matching loss recovers the gradient of the diffusion model’s evidence lower bound, so that optimizing it yields the maximum likelihood estimate of the data distribution song2021maximum . However, optimizing Equation 7 is costly, as it requires computing the gradient through the flow model. As shown in wang2023prolificdreamer this can be circumvented by reformulating the regularizer in terms of the Wasserstein gradient flow:
$\nabla_{\mu_{x}}\mathcal{R}(x)=\mathbb{E}_{q(x_{t}|y)}\left[\omega(t)(\underbrace{\nabla_{x}\log q(x_{t}|y)}_{\text{score of noisy variational distribution}}-\underbrace{\nabla_{x}\log p(x_{t})}_{\text{score of noisy prior distribution}})\right]$ (8)
Note that optimizing only the regularization term, without the data term, at test time is equivalent to the objective of Score Distillation Sampling (SDS) pooledreamfusion .
4 Method
Flow Formulation. The variational formulation in Equation 7 is formulated for the score, but can be reformulated into a denoising or $\epsilon_{\theta}$ parametrization song2021scorebased ; mardani2024a . However, we are interested in a variational objective that depends on the velocity field $v_{\theta}(x_{t},t)$, which characterizes the probabilistic trajectory that connects the noise and data distributions.
Proposition 1.
We propose to replace the score-based regularizer in the standard variational objective with a flow matching formulation, resulting in the following objective function:
$\arg\min_{q(x_{0}|y)}\underbrace{\mathbb{E}_{q(x_{0}|y)}\left[\frac{\|y-f(\mu_{x})\|^{2}}{2\sigma_{\nu}^{2}}\right]}_{\text{$\mathcal{D}(x,y)$}}+\underbrace{\int_{0}^{T}\lambda_{\mathcal{R}}(t)~{}\mathbb{E}_{q(x_{t}|y)}\left[\left\|v_{\theta}(x_{t},t)-u_{t}(x_{t}|\epsilon)\right\|^{2}\right]dt}_{\text{$\mathcal{R}(x)$}}$ (9)
The flow-matching term that defines the regularizer arises by re-parameterizing the variational distribution to $q(x_{t}|y)=\mathcal{N}((1-t)\mu_{x},t^{2}\,I)$. This corresponds to sampling via the deterministic map $\psi_{t}(x_{0}\mid\epsilon)=(1-t)\,x_{0}+t\,\epsilon$, with $\epsilon\sim\mathcal{N}(0,I)$. By re-formulating the score in terms of the target velocity field $u_{t}$, we get:
$\nabla_{x}\log q(x_{t}|y)=-\frac{(1-t)u_{t}(x_{t}|\epsilon)+x_{t}}{t}$ (10)
For the learned velocity $v_{\theta}(x_{t},t)$ a similar approximation holds – for a full derivation please see the supplementary material subsection A.2.
$v_{\theta}(x_{t},t)\approx\frac{-t\nabla_{x}\log p(x_{t})-x_{t}}{1-t}$ (11)
We can therefore approximate the score of the noisy prior with our learned velocity field $v_{\theta}$
$\nabla_{x}\log p(x_{t})\approx-\frac{(1-t)v_{\theta}(x_{t},t)+x_{t}}{t}.$ (12)
Hard Data Consistency. Existing variational posterior sampling approaches mardani2024a ; zilberstein2025repulsive impose soft constraints on the data fidelity term $\mathcal{D}(x,y)$. In contrast, recent work song2024solving has demonstrated that, when sampling from latent diffusion models, enforcing hard data consistency generally leads to better reconstructions with improved visual fidelity. Our method shares this motivation, but differs in that we optimize over a variational distribution, i.e., we compute $\min~{}\mathbb{E}_{q(x_{0}|y)}[-\log p(y|x_{0})]$. An additional advantage of this variational setup is that it allows us to initialize the optimization variable with an adjoint based initialization $\mu_{x}=E(A^{\top}y)$, with $E$ being the encoder of the VAE.
Accuracy Calibration. As our framework evaluates the trajectory at each timestep, we aim to weight the regularizer’s contribution according to its reliability. It has been shown that the difficulty of the prediction task depends on the network parameterization, as well as on the specific timestep $t$ Karras2022edm . Since the regularization term $\mathcal{R}(x)$ in our approach is equivalent to the training objective of the pre-trained flow model, we can easily weight it by the expected model error, which we calibrate on a small set of images. Specifically, we sample $N$ calibration images and compute the conditional flow matching objective for 100 linearly spaced time steps between 0 and 1, then average the error over all images to obtain the expected model error at each time step. Different functions of the model error can be chosen as weight for the regularizer. We choose:
$\lambda_{\mathcal{R}}(t)=\frac{1}{N}\left(\sum_{i=1}^{N}\left\|v_{\theta}(x^{(i)}_{t},t)-u_{t}(x^{(i)}_{t}\mid\epsilon)\right\|^{2}\right)^{-1}$ (13)
and set $\lambda_{\mathcal{R}}(t)=0$ for all $t<0.2$, since the accuracy of SD3 is heavily degraded for low noise levels.
Deterministic Trajectory Adjustment. Score distillation sampling relies on the assumption that $x_{t}=(1-t)\mu_{x}+t\epsilon$ lies in a region of the learned prior that has reasonably high support/density. In practice, this is not always the case. When not tightly conditioned (usually with extensive text prompts), even the best available diffusion models assign low density to many plausible regions of the latent space, leading to bad gradient steps. Therefore, we increase the probability of $p(x_{t})$ by additionally conditioning $x_{t}$ on the estimated "end-point" $\mu$.
Proposition 2.
We introduce a reparameterized variational distribution with a mean that linearly interpolates between the posterior mean $\mu_{x}$ and a model-guided  $\hat{x}_{1}$:
$q(x_{t}\mid y)=\mathcal{N}\left((1-t)\mu_{x}+t\alpha\hat{x}_{1},\;t^{2}(1-\alpha^{2})I\right),$ (14)
where $\hat{x}_{1}=x_{t+\delta t}+(1-t-\delta t)v_{\theta}(x_{t+\delta t},t+\delta t)$ is a single-step velocity-based predictor, and $\alpha\in[0,1]$ controls the trade-off between deterministic guidance and random noise. This reparameterization induces the following reference velocity field:
$u_{t}(x_{t}\mid\epsilon)=\frac{\alpha\hat{x}_{1}+\sqrt{1-\alpha^{2}}\epsilon-x_{t}}{1-t}.$ (15)
Intuitively, changing the formulation in this manner ensures that the model relocates the sample to its expected position on the learned manifold, rather than injecting arbitrary noise, which could drive it in a direction that has high prior likelihood but is not consistent with the observation. To further encourage exploration and avoid collapsing onto the trajectory of the adjoint measurement, we inject an additional, stochastic component $\epsilon$ during this process. A full derivation can be found in the supplementary material, subsection A.2.
4.1 Algorithm
The following pseudo-code summarizes our method, integrating all the components discussed above. We adapt the standard scheme mardani2024a ; zilberstein2025repulsive of linearly traversing time in a descending manner and stop at $t=0.2$ as explained in section 4. We choose $\alpha=1-t$. Gradient updates to enforce hard data consistency are performed using stocastic gradient descent. For further implementation details and ablations, see subsection A.4
Input: $\mu_{x}=A^{\top}(y)$, $\lambda_{R}$, $\alpha$, $y$, $A$, $v_{\theta}$
Output: $\mu_{x}$
$\hat{\epsilon}\sim\mathcal{N}(0,I)$;
  $\triangleright$ initial noise sample
for $t\leftarrow 1$ to 00 by $-\Delta t$ do
      $x_{t}\leftarrow(1-t)\,\mu_{x}+t\,\hat{\epsilon}$;
        $\triangleright$ sample noisy latent
       $u_{t}(x_{t}\mid\hat{\epsilon})\leftarrow\dfrac{\hat{\epsilon}-x_{t}}{1-t}$;
      $\nabla_{\mu_{x}}R\leftarrow v_{\theta}(x_{t},t)-u_{t}(x_{t}\mid\hat{\epsilon})$;
      $\mu_{x}\leftarrow\mu_{x}-\lambda_{R}\,\nabla_{\mu_{x}}R$;
        $\triangleright$ update w.r.t. regularizer
      $\mu_{x}\leftarrow\arg\min_{\mu_{x}}\|y-A(\mu_{x})\|^{2}$;
        $\triangleright$ hard data consistency
      $\epsilon\sim\mathcal{N}(0,I)$;
      $\hat{x}_{1}\leftarrow x_{t}+(1-t)\,v_{\theta}(x_{t},t)$;
        $\triangleright$ predict deterministic noise
       $\hat{\epsilon}\leftarrow\alpha\,\hat{x}_{1}+\sqrt{1-\alpha^{2}}\,\epsilon$;
        $\triangleright$ update noise estimate
Algorithm 1 The FLAIR solver for inverse imaging problems
5 Experiments
We evaluate the performance of FLAIR on a variety of inverse imaging tasks and compare against several baselines, using the SD3 backbone without any fine-tuning. We used several metrics including SSIM wang2004image , LPIPS zhang2018unreasonable and patchwise FID heusel2017gans (pFID) to comprehensively assess the perceptual and quantitative quality of the reconstructions. FID is computed using InceptionV3 features on patches of 256x256 resolution. All experiments were performed on a single NVidia RTX 4090 GPU with 24GB of VRAM. For completeness we also show PSNR values, but point out that the metric is not well suited for our purposes: PSNR favors the posterior mean, while the goal of the variational approach is to sample from the posterior distribution. Accordingly, PSNR is known to prefer over-smoothed, blurry outputs over sharp ones with even the slightest misalignment of high frequency content blau2018perception .
5.1 Setup
Datasets. We utilize two high-resolution image datasets: FFHQ karras2019style and DIV2K Agustsson_2017_CVPR_Workshops . FFHQ consists of 70k diverse face images at 1024×1024 resolution of which we take the first 1000 samples. It is covering variations in age, pose, lighting, and ethnicity. DIV2K contains 800 high-quality images in 2K resolution that span a range of natural scenes with varied textures and structures.
Baselines. Our method is benchmarked against several recent inverse imaging solvers based on posterior sampling. Specifically, we compare to ReSample song2024solving , FlowDPS kim2025flowdps , FlowChef patel2024flowchef , and RSD zilberstein2025repulsive . The latter is used without repulsive term as it delivers better results. To ensure a fair and meaningful comparison, all methods are evaluated with the same number of function evaluations.
Problem Setting. We run and evaluate all methods at a fixed output resolution of 768$\times$768 pixels. For single image super-resolution, we consider scaling factors of 8×8$\times$8 × and 12×12$\times$12 ×. The corresponding low-resolution inputs are generated by bicubic downsampling. Motion blur is simulated with a blur kernel of size 61. For box inpainting, we mask large, continuous rectangles that cover approximately one third of the observation. All synthesized observations are corrupted with additive Gaussian noise, with standard deviation $\sigma_{\nu}$ of 0.5%. For inference on the FFHQ dataset, we use a predefined text prompt of the form "A high quality photo of a face", and for DIV2k "A high quality photo of" concatenated with an image-specific description retrieved by applying DAPE Wu_2024_CVPR to the observation.
5.2 Experimental Results
Inverse Problems. Our experiments clearly demonstrate that FLAIR outperforms existing flow-based approaches in terms of all perceptual metrics, see Table 1.
In the case of image inpainting, our method produces high-quality reconstructions that fully leverage the power of the generative model and blend naturally into the surrounding context, avoiding degradations and artifacts that we observe in the baselines. In particular, FlowDPS tends to produce implausible textures in the inpainted regions, while FlowChef regularly fails to generate semantically consistent content at all. For single-image super-resolution, FLAIR consistently delivers the most perceptually convincing and realistic outputs. Notably, the FID scores remain low for both ×$\times 8$ and ×$\times 12$ magnification, indicating an effective usage of the generative prior to overcome the increasing ill-posedness. Again, FlowDPS suffers from blur and low texture quality, whereas FlowChef tends to lose semantic coherence. In motion deblurring, FLAIR also restores sharper and semantically more credible content than competing approaches, which often suffer from residual blur or inconsistent details. The boost in reconstruction quality is quantitatively reflected by all metrics, confirming that FLAIR reconstructs images with high fidelity. For further qualitative examples, please see subsection A.9.
Figure 2: Qualitative comparison. FLAIR produces posterior samples of high perceptual quality while maintaining high data likelihood. Best viewed zoomed in.
Table 1: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$.
SR $\times$8 SR $\times$12 Motion Deblurring Inpainting
[…]
FFHQ 768$\times$768
ReSample 0.400 331 109 762 191 80.40033110976219180.400\,331\,109\,762\,191\,80.400 331 109 762 191 8 55.586 227 416 992 1955.5862274169921955.586\,227\,416\,992\,1955.586 227 416 992 19 0.814 689 397 811 889 60.81468939781188960.814\,689\,397\,811\,889\,60.814 689 397 811 889 6 26.367 465 972 900 3926.3674659729003926.367\,465\,972\,900\,3926.367 465 972 900 39 0.473 518 252 372 741 70.47351825237274170.473\,518\,252\,372\,741\,70.473 518 252 372 741 7 80.268 859 863 281 2580.2688598632812580.268\,859\,863\,281\,2580.268 859 863 281 25 0.785 562 992 095 947 30.78556299209594730.785\,562\,992\,095\,947\,30.785 562 992 095 947 3 25.469 425 201 416 01625.46942520141601625.469\,425\,201\,416\,01625.469 425 201 416 016 0.457 411 199 808 120 70.45741119980812070.457\,411\,199\,808\,120\,70.457 411 199 808 120 7 82.892 143 249 511 7282.8921432495117282.892\,143\,249\,511\,7282.892 143 249 511 72 0.787 788 689 136 505 10.78778868913650510.787\,788\,689\,136\,505\,10.787 788 689 136 505 1 25.446 083 068 847 65625.44608306884765625.446\,083\,068\,847\,65625.446 083 068 847 656 0.365 910 947 322 845 460.365910947322845460.365\,910\,947\,322\,845\,460.365 910 947 322 845 46 70.832 069 396 972 6670.8320693969726670.832\,069\,396\,972\,6670.832 069 396 972 66 0.827 21.827 827 453 613 2821.8278274536132821.827\,827\,453\,613\,2821.827 827 453 613 28
FlowDPS 0.374 372 541 904 449 460.374372541904449460.374\,372\,541\,904\,449\,460.374 372 541 904 449 46 40.299 484 252 929 6940.2994842529296940.299\,484\,252\,929\,6940.299 484 252 929 69 0.756 326 615 810 394 30.75632661581039430.756\,326\,615\,810\,394\,30.756 326 615 810 394 3 29.238 021 850 585 93829.23802185058593829.238\,021\,850\,585\,93829.238 021 850 585 938 0.413 043 200 969 696 040.413043200969696040.413\,043\,200\,969\,696\,040.413 043 200 969 696 04 47.244 354 248 046 87547.24435424804687547.244\,354\,248\,046\,87547.244 354 248 046 875 0.741 228 818 893 432 60.74122881889343260.741\,228\,818\,893\,432\,60.741 228 818 893 432 6 28.05 0.431 281 954 050 064 10.43128195405006410.431\,281\,954\,050\,064\,10.431 281 954 050 064 1 49.485 298 156 738 2849.4852981567382849.485\,298\,156\,738\,2849.485 298 156 738 28 0.731 617 748 737 335 20.73161774873733520.731\,617\,748\,737\,335\,20.731 617 748 737 335 2 27.637 281 417 846 6827.6372814178466827.637\,281\,417\,846\,6827.637 281 417 846 68 0.344 38.3 0.770 720 958 709 716 80.77072095870971680.770\,720\,958\,709\,716\,80.770 720 958 709 716 8 19.191 648 483 276 36719.19164848327636719.191\,648\,483\,276\,36719.191 648 483 276 367
RSD 0.391 473 561 525 344 850.391473561525344850.391\,473\,561\,525\,344\,850.391 473 561 525 344 85 51.667 427 062 988 2851.6674270629882851.667\,427\,062\,988\,2851.667 427 062 988 28 0.776 033 043 861 389 20.77603304386138920.776\,033\,043\,861\,389\,20.776 033 043 861 389 2 29.690 019 607 543 94529.69001960754394529.690\,019\,607\,543\,94529.690 019 607 543 945 0.462 387 174 367 904 660.462387174367904660.462\,387\,174\,367\,904\,660.462 387 174 367 904 66 71.734 550 476 074 2271.7345504760742271.734\,550\,476\,074\,2271.734 550 476 074 22 0.743 28.108 963 012 695 31228.10896301269531228.108\,963\,012\,695\,31228.108 963 012 695 312 0.457 823 425 531 387 330.457823425531387330.457\,823\,425\,531\,387\,330.457 823 425 531 387 33 77.319 137 573 242 1977.3191375732421977.319\,137\,573\,242\,1977.319 137 573 242 19 0.743 024 885 654 449 50.74302488565444950.743\,024\,885\,654\,449\,50.743 024 885 654 449 5 27.67 0.478 281 795 978 546 140.478281795978546140.478\,281\,795\,978\,546\,140.478 281 795 978 546 14 73.293 403 625 488 2873.2934036254882873.293\,403\,625\,488\,2873.293 403 625 488 28 0.735 569 655 895 233 20.73556965589523320.735\,569\,655\,895\,233\,20.735 569 655 895 233 2 21.97
FlowChef 0.341 30.5 0.760 072 886 943 817 10.76007288694381710.760\,072\,886\,943\,817\,10.760 072 886 943 817 1 28.424 997 329 711 91428.42499732971191428.424\,997\,329\,711\,91428.424 997 329 711 914 0.373 41.0 0.730 421 423 912 048 30.73042142391204830.730\,421\,423\,912\,048\,30.730 421 423 912 048 3 26.996 358 871 459 9626.9963588714599626.996\,358\,871\,459\,9626.996 358 871 459 96 0.406 35.0 0.716 119 468 212 127 70.71611946821212770.716\,119\,468\,212\,127\,70.716 119 468 212 127 7 25.809 129 714 965 8225.8091297149658225.809\,129\,714\,965\,8225.809 129 714 965 82 0.394 163 548 946 380 60.39416354894638060.394\,163\,548\,946\,380\,60.394 163 548 946 380 6 71.046 714 782 714 8471.0467147827148471.046\,714\,782\,714\,8471.046 714 782 714 84 0.780 103 147 029 876 70.78010314702987670.780\,103\,147\,029\,876\,70.780 103 147 029 876 7 18.184 221 267 700 19518.18422126770019518.184\,221\,267\,700\,19518.184 221 267 700 195
Ours 0.213 235 303 759 574 90.21323530375957490.213\,235\,303\,759\,574\,90.213 235 303 759 574 9 13.266 211 509 704 5913.2662115097045913.266\,211\,509\,704\,5913.266 211 509 704 59 0.777 29.54 0.270 564 854 145 050 050.270564854145050050.270\,564\,854\,145\,050\,050.270 564 854 145 050 05 16.201 229 095 458 98416.20122909545898416.201\,229\,095\,458\,98416.201 229 095 458 984 0.740 209 460 258 483 90.74020946025848390.740\,209\,460\,258\,483\,90.740 209 460 258 483 9 27.714 468 002 319 33627.71446800231933627.714\,468\,002\,319\,33627.714 468 002 319 336 0.236 155 018 210 411 070.236155018210411070.236\,155\,018\,210\,411\,070.236 155 018 210 411 07 10.732 836 723 327 63710.73283672332763710.732\,836\,723\,327\,63710.732 836 723 327 637 0.772 29.611 394 882 202 1529.6113948822021529.611\,394\,882\,202\,1529.611 394 882 202 15 0.183 834 627 270 698 550.183834627270698550.183\,834\,627\,270\,698\,550.183 834 627 270 698 55 8.741 712 570 190 438.741712570190438.741\,712\,570\,190\,438.741 712 570 190 43 0.827 794 373 035 430 90.82779437303543090.827\,794\,373\,035\,430\,90.827 794 373 035 430 9 23.693 054 199 218 7523.6930541992187523.693\,054\,199\,218\,7523.693 054 199 218 75
DIV2K 768$\times$768
ReSample $0.533$ 54.997 371 673 583 9854.9973716735839854.997\,371\,673\,583\,9854.997 371 673 583 98 $0.625$ $22.34$ 0.642 829 596 996 307 40.64282959699630740.642\,829\,596\,996\,307\,40.642 829 596 996 307 4 88.051 658 630 371 188.051658630371188.051\,658\,630\,371\,188.051 658 630 371 1 0.561 500 668 525 695 80.56150066852569580.561\,500\,668\,525\,695\,80.561 500 668 525 695 8 $20.85$ 0.555 853 188 037 872 30.55585318803787230.555\,853\,188\,037\,872\,30.555 853 188 037 872 3 79.654 029 846 191 479.654029846191479.654\,029\,846\,191\,479.654 029 846 191 4 0.617 21.792 427 062 988 2821.7924270629882821.792\,427\,062\,988\,2821.792 427 062 988 28 0.285 51.859 077 453 613 2851.8590774536132851.859\,077\,453\,613\,2851.859 077 453 613 28 0.796 22.676 401 138 305 66422.67640113830566422.676\,401\,138\,305\,66422.676 401 138 305 664
FlowDPS 0.476 54.160 160 064 697 26654.16016006469726654.160\,160\,064\,697\,26654.160 160 064 697 266 0.567 384 541 034 698 50.56738454103469850.567\,384\,541\,034\,698\,50.567 384 541 034 698 5 23.005 729 675 292 9723.0057296752929723.005\,729\,675\,292\,9723.005 729 675 292 97 0.547 238 528 728 485 10.54723852872848510.547\,238\,528\,728\,485\,10.547 238 528 728 485 1 67.424 385 070 800 7867.4243850708007867.424\,385\,070\,800\,7867.424 385 070 800 78 0.528 21.79 0.558 460 950 851 440 40.55846095085144040.558\,460\,950\,851\,440\,40.558 460 950 851 440 4 60.432 220 458 984 37560.43222045898437560.432\,220\,458\,984\,37560.432 220 458 984 375 0.535 993 039 608 001 70.53599303960800170.535\,993\,039\,608\,001\,70.535 993 039 608 001 7 21.877 470 016 479 49221.87747001647949221.877\,470\,016\,479\,49221.877 470 016 479 492 0.327 635 318 040 847 80.32763531804084780.327\,635\,318\,040\,847\,80.327 635 318 040 847 8 34.2 0.691 909 790 039 062 50.69190979003906250.691\,909\,790\,039\,062\,50.691 909 790 039 062 5 21.714 540 481 567 38321.71454048156738321.714\,540\,481\,567\,38321.714 540 481 567 383
RSD 0.538 948 655 128 4790.5389486551284790.538\,948\,655\,128\,4790.538 948 655 128 479 60.930 324 554 443 3660.9303245544433660.930\,324\,554\,443\,3660.930 324 554 443 36 0.591 312 229 633 331 30.59131222963333130.591\,312\,229\,633\,331\,30.591 312 229 633 331 3 23.445 989 608 764 6523.4459896087646523.445\,989\,608\,764\,6523.445 989 608 764 65 0.683 576 822 280 8830.6835768222808830.683\,576\,822\,280\,8830.683 576 822 280 883 95.726 249 694 824 2295.7262496948242295.726\,249\,694\,824\,2295.726 249 694 824 22 0.523 275 852 203 369 10.52327585220336910.523\,275\,852\,203\,369\,10.523 275 852 203 369 1 21.964 975 357 055 66421.96497535705566421.964\,975\,357\,055\,66421.964 975 357 055 664 0.637 668 132 781 982 40.63766813278198240.637\,668\,132\,781\,982\,40.637 668 132 781 982 4 97.646 888 732 910 1697.6468887329101697.646\,888\,732\,910\,1697.646 888 732 910 16 0.551 110 327 243 804 90.55111032724380490.551\,110\,327\,243\,804\,90.551 110 327 243 804 9 22.10 0.463 778 406 381 607 060.463778406381607060.463\,778\,406\,381\,607\,060.463 778 406 381 607 06 63.870 880 126 953 12563.87088012695312563.870\,880\,126\,953\,12563.870 880 126 953 125 0.677 523 434 162 139 90.67752343416213990.677\,523\,434\,162\,139\,90.677 523 434 162 139 9 23.23
FlowChef 0.490 322 709 083 557 130.490322709083557130.490\,322\,709\,083\,557\,130.490 322 709 083 557 13 48.6 0.539 492 428 302 764 90.53949242830276490.539\,492\,428\,302\,764\,90.539 492 428 302 764 9 21.842 645 645 141 621.842645645141621.842\,645\,645\,141\,621.842 645 645 141 6 0.525 55.3 0.492 154 121 398 925 80.49215412139892580.492\,154\,121\,398\,925\,80.492 154 121 398 925 8 20.521 999 359 130 8620.5219993591308620.521\,999\,359\,130\,8620.521 999 359 130 86 0.561 51.3 0.486 047 297 716 140 750.486047297716140750.486\,047\,297\,716\,140\,750.486 047 297 716 140 75 19.899 375 915 527 34419.89937591552734419.899\,375\,915\,527\,34419.899 375 915 527 344 0.488 857 418 298 721 30.48885741829872130.488\,857\,418\,298\,721\,30.488 857 418 298 721 3 86.286 514 282 226 5686.2865142822265686.286\,514\,282\,226\,5686.286 514 282 226 56 0.659 262 299 537 658 70.65926229953765870.659\,262\,299\,537\,658\,70.659 262 299 537 658 7 20.874 286 651 611 32820.87428665161132820.874\,286\,651\,611\,32820.874 286 651 611 328
Ours 0.352 998 495 101 928 70.35299849510192870.352\,998\,495\,101\,928\,70.352 998 495 101 928 7 26.456 972 122 192 38326.45697212219238326.456\,972\,122\,192\,38326.456 972 122 192 383 0.607 23.30 0.421 490 550 041 198 730.421490550041198730.421\,490\,550\,041\,198\,730.421 490 550 041 198 73 32.118 663 787 841 832.118663787841832.118\,663\,787\,841\,832.118 663 787 841 8 0.525 365 948 677 0630.5253659486770630.525\,365\,948\,677\,0630.525 365 948 677 063 21.385 475 158 691 40621.38547515869140621.385\,475\,158\,691\,40621.385 475 158 691 406 0.314 553 380 012 512 20.31455338001251220.314\,553\,380\,012\,512\,20.314 553 380 012 512 2 21.111 848 831 176 75821.11184883117675821.111\,848\,831\,176\,75821.111 848 831 176 758 0.652 780 890 464 782 70.65278089046478270.652\,780\,890\,464\,782\,70.652 780 890 464 782 7 24.444 496 154 785 15624.44449615478515624.444\,496\,154\,785\,15624.444 496 154 785 156 0.162 881 195 545 196 530.162881195545196530.162\,881\,195\,545\,196\,530.162 881 195 545 196 53 11.026 673 316 955 56611.02667331695556611.026\,673\,316\,955\,56611.026 673 316 955 566 0.814 652 204 513 549 80.81465220451354980.814\,652\,204\,513\,549\,80.814 652 204 513 549 8 23.752 101 898 193 3623.7521018981933623.752\,101\,898\,193\,3623.752 101 898 193 36
Posterior Variance. To demonstrate that FLAIR does not suffer from mode collapse, we assess the posterior variance $\operatorname*{Var}[x|z]$ for the task of ×$\times 12$ Super Resolution, by drawing 32 samples for a fixed observation $z$ and computing their pixel-wise variance. We conduct that experiment for our FLAIR approach, for RSD with repulsive term, and for FlowDPS kim2025flowdps . The example in Figure 3 illustrates that FLAIR has the highest sample diversity, which is also reflected in the corresponding variance maps. Notably, the sample variance is concentrated in regions with high-frequency textures. This indicates that our method reliably reconstructs the posterior, whose low-frequency part is, in the super-resolution setting, tightly constrained by the likelihood term.
Figure 3: Zoomed-in reconstructions for x12 Super Resolution. We show posterior samples (col. 1–4) of FLAIR, FlowDPS, and RSD, posterior mean and standard deviation (over 32 samples, col. 5,6). 00 Refer to caption $0.16$
Editing. Beyond image restoration, we observe that our method also performs remarkably well for text-based image editing, by simply presenting suitable target prompts during inpainting. Figure 4 illustrates a variety of edited images generated from the same photograph with the help of the depicted masks and prompts.
Figure 4: Edited images shown alongside original, with prompts: "A high resolution portrait of a…"
5.3 Ablation Studies
We systematically analyze the impact of key design choices in our method. Specifically, we ablate the deterministic trajectory adjustment, the use of hard data consistency, and the calibration of the regularizer weight for ×$\times 12$ super-resolution, using a subset of 100 samples from the FFHQ and DIV2K datasets. Quantitative and qualitative results are shown in Table 2 and Figure 5, respectively.
Hard Data Consistency (HDC). Dropping the hard data consistency degrades both metrics, with PSNR being particularly affected due to poorer alignment with the input observation, which is also evident in the visual example: the reconstruction is plausible, but deviates from the observation.
Deterministic Trajectory Adjustment (DTA). The biggest performance drop compared to the full setup occurs when removing the deterministic trajectory adjustment, as standard noise injection harms the gradient updates in low-density regions of the prior. The reconstruction appears overly smooth and lacks texture details.
Calibrated Regularizer Weight (CRW). Replacing our calibrated regularizer weight with $\lambda_{\mathcal{R}}(t)=t$ also has a strong impact on perceptual quality: the result is visibly blurred if one ignores the changing accuracy of the regularizer along the flow trajectory.
Figure 5: Qualitative samples from the ablation study on $\times$12 Super Resolution.
Table 2: Ablation study for $\times$12 super-resolution on DIV2K and FFHQ. Model components are individually switched on or off.
HDC DTA CRW FFHQ DIV2K
LPIPS ↓ PSNR ↑ LPIPS ↓ PSNR ↑
0.259 27.45 0.427 21.05
0.297 27.17 0.467 20.82
0.432 27.20 0.622 21.69
0.363 28.58 0.583 21.98
0.392 28.33 0.605 21.99
Legend. HDC: Hard Data Consistency; DTA: Deterministic Trajectory Adjustment;
CRW: Calibrated Regularizer Weight. ✓ = included, ✗ = ablated.
6 Conclusion and Limitations
We have presented FLAIR, a training-free variational framework for inverse problems that uses a flow-based generative model as its image prior. By combining the power of (latent) flow-based models with a principled reconstruction of the posterior distribution, FLAIR addresses key limitations of existing methods. First, it is able to recover atypical modes, by combining the degradation-agnostic flow matching loss with deterministic trajectory adjustments. Second, it enables hard data consistency without sacrificing sample diversity, by decoupling the data consistency constraint from the regularization, while adaptively reweighting the latter according to its expected accuracy, calibrated offline. Experiments with different image datasets and tasks confirm that FLAIR consistently achieves higher reconstruction quality than existing baselines based on either flow matching or denoising diffusion. Notably, our proposed method achieves, at the same time, excellent perceptual quality, close adherence to the input observations, and high sample diversity.
Evidently, FLAIR inherits the limitations of the underlying generative model, in our implementation Stable Diffusion 3. These include biases caused by the selection of training data, constraints w.r.t. the output resolution, and a limited ability to recover out-of-distribution modes. Furthermore, our approach introduces additional hyper-parameters needed to control the deterministic trajectory adjustment. We note that high fidelity image restoration methods can potentially be misused for unethical image manipulations.
Supplementary Material
In the following, we provide detailed line-by-line derivations of the mathematical formulations used in the paper, as well as additional implementation details and experimental results.
Appendix A Derivations
A.1 Derivation of flow-based variational formulation
The linear conditional flow and it’s corresponding velocity are defined as:
$\psi_{t}(x_{0}\mid\epsilon)=(1-t)\,x_{0}+t\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)\;,$ (16)
$u_{t}(x_{t}|\epsilon)=\frac{\operatorname{d}\!{\psi_{t}}}{\operatorname{d}\!{t}}(\psi_{t}^{-1}(x_{t}|\epsilon)|\epsilon)\;.$ (17)
The score of the noisy variational distribution can be analytically computed with:
$q(x_{t}|y)=\mathcal{N}((1-t)\mu_{x},t^{2}\,I)\;,$ (18)
$\nabla_{x_{t}}\log q(x_{t}|y)=-\frac{\epsilon}{t}\;.$ (19)
We compute $\frac{\operatorname{d}\!{\psi_{t}}}{\operatorname{d}\!{t}}(x_{0}|\epsilon)=-x_{0}+\epsilon$ and $\psi_{t}^{-1}(x_{t}|\epsilon)$ and insert it into Equation 17:
$u_{t}(x_{t}|\epsilon)=\frac{\epsilon-x_{t}}{1-t}\;.$ (20)
Solving Equation 20 for $\epsilon$ and inserting in Equation 19 gives:
$\nabla_{x_{t}}\log q(x_{t}|y)=-\frac{(1-t)u_{t}(x_{t}|\epsilon)+x_{t}}{t}\;.$ (21)
For the learned velocity $v_{\theta}(x_{t},t)$ a similar approximation holds:
$v_{\theta}(x_{t},t)\approx\frac{-t\nabla_{x}\log p(x_{t})-x_{t}}{1-t}\;.$ (22)
Hence, we can approximate the score of the noisy prior with our learned velocity field $v_{\theta}$
$\nabla_{x_{t}}\log p(x_{t})\approx-\frac{(1-t)v_{\theta}(x_{t},t)+x_{t}}{t}\;,$ (23)
and we see that for $\omega(t)=\frac{t}{1-t}$ we obtain the conditional flow matching objective for $\mathcal{R}(x)$. We therefore set $\omega(t)=\frac{t}{1-t}$ and end up at our final objective:
$\arg\min_{q(x_{0}|y)}\underbrace{\mathbb{E}_{q(x_{0}|y)}\left[\frac{\|y-f(\mu_{x})\|^{2}}{2\nu^{2}}\right]}_{\text{$\mathcal{D}(x,y)$}}+\underbrace{\int_{0}^{T}~{}\mathbb{E}_{q(x_{t}|y)}\left[\left\|v_{\theta}(x_{t},t)-u_{t}(x_{t}|\epsilon)\right\|^{2}\right]dt}_{\text{$\mathcal{R}(x)$}}\;.$ (24)
Again, the gradient step for the regularizer becomes:
$\nabla_{\mu_{x}}\mathcal{R}(x)=\mathbb{E}_{q(x_{t}|y)}\left[v_{\theta}(x_{t},t)-u_{t}(x_{t}|\epsilon)\right]\;.$ (25)
A.2 Derivation of trajectory adjusted flow-based variational formulation
To achieve the proposed trajectory adjustment, we modify the forward process to:
$\hat{x}_{1}=x_{t+dt}+(1-t-dt)v_{\theta}(x_{t+dt},t+dt)\;,$ (26)
$x_{t}=(1-t)\mu_{x}+t\underbrace{(\alpha\hat{x}_{1}+\sqrt{1-\alpha^{2}}\epsilon)}_{\text{$\hat{\epsilon}$}}\;,$ (27)
where $\hat{x}_{1}$ is the noise vector prediction from the last optimization iteration. This induces a variational distribution:
$q(x_{t}\mid y)=\mathcal{N}\left((1-t)\mu_{x}+t\alpha\hat{x}_{1},\,t^{2}(1-\alpha^{2})I\right)\;,$ (28)
leading to a score of
$\nabla_{x_{t}}\log q(x_{t}\mid y)=-\frac{1}{t^{2}(1-\alpha^{2})}\cdot t\sqrt{1-\alpha^{2}}\epsilon=-\frac{\epsilon}{t\sqrt{1-\alpha^{2}}}\;.$ (29)
The velocity field is again computed by Equation 17. We start by defining the flow:
$\psi_{t}(x_{0}\mid\epsilon)=(1-t)x_{0}+t\left(\alpha\hat{x}_{1}+\sqrt{1-\alpha^{2}}\epsilon\right)\;.$ (30)
The resulting derivative reads
$\frac{d}{dt}\psi_{t}(x_{0}\mid\epsilon)=\alpha\hat{x}_{1}-x_{0}+\sqrt{1-\alpha^{2}}\epsilon\;,$ (31)
and the inverse becomes
$x_{0}=\psi_{t}^{-1}(x_{t}\mid\epsilon)=\frac{x_{t}-t\alpha\hat{x}_{1}-t\sqrt{1-\alpha^{2}}\epsilon}{1-t}\;.$ (32)
Plugging these results into Equation 17:
$u_{t}(x_{t}\mid\epsilon)=\frac{\alpha\hat{x}_{1}+\sqrt{1-\alpha^{2}}\epsilon-x_{t}}{1-t}\;.$ (33)
A.3 Derivation of Score from Flow
The score matching objective reads as:
[…]
where,
$-\frac{1}{\sigma(t)^{2}}\left(x_{t}-\mu(x_{0},t)\right)=\nabla_{x_{t}}\log p_{t}(x_{t}\mid x_{0}),$ (35)
with $p_{t}(x_{t}\mid x_{0})=\mathcal{N}(\mu_{t}(x_{0}),\sigma_{t}^{2}I)$. Note that as usual we assume $\mu_{t}(x_{0})$ being linear in $x_{0}$.
Equation 34 is solved by:
$\nabla_{x_{t}}\log p_{t}(x_{t})=\mathbb{E}_{p_{t}(x_{0}|x_{t})}\left[\nabla_{x_{t}}\log p_{t}(x_{t}|x_{0})\right],$ (36)
and can be written as:
$\nabla_{x_{t}}\log p_{t}(x_{t})=\frac{-(x_{t}-\mu(\mathbb{E}[x_{0}\mid x_{t}],t))}{\sigma(t)^{2}}.$ (37)
In the case of OT flow-matching, we obtain
$x_{t}=(1-t)x_{0}+tx_{1},$ (38)
$x_{1}\sim\mathcal{N}(0,\mathrm{Id})$ and $p(x_{t}|x_{0})=\mathcal{N}((1-t)x_{0},t^{2})$. The optimal velocity under the flow matching loss is given by:
$v^{*}(x_{t},t)=\mathbb{E}[x_{1}-x_{0}\mid x_{t}].$ (39)
Expressing $x_{1}=\frac{x_{t}-(1-t)x_{0}}{t}$, we can insert into Equation 39 and obtain:
$\mathbb{E}[x_{0}\mid x_{t}]=x_{t}-t\mathbb{E}[x_{1}-x_{0}\mid x_{t}].$ (40)
Inserting in Equation 37 leads to:
$\nabla_{x_{t}}\log p_{t}(x_{t})=-\frac{x_{t}-(1-t)(x_{t}-t\mathbb{E}[x_{1}-x_{0}\mid x_{t}])}{t^{2}},$ (41)
which for $v^{*}(x_{t},t)=\mathbb{E}[x_{1}-x_{0}\mid x_{t}]$ reads as:
$\nabla_{x_{t}}\log p(x_{t})\approx-\frac{(1-t)v_{\theta}(x_{t},t)+x_{t}}{t}.$ (42)
A.4 Implementation details
Flow Model and Regularizer Settings. As flow matching model, we us Stable Diffusion 3.5-Medium, which has been released under the Stability Community License. The classifier-free guidance scale is set to 2 for all experiments. To minimize the regularization term, we use stochastic gradient descent with a learning rate of $1$.
Data Likelihood Term. We use stochastic gradient descent for the minimization of the data term towards hard data consistency. For numerical stability, the squared error is summed over all measurements instead of computing the mean. The learning rate has to be adjusted accordingly, to compensate for the varying number of measurements $y$. Moreover, the minimization is terminated with early stopping once the likelihood term reaches $1\times 10^{-4}\cdot\text{len}(y)$, to not overfit the noise in the image observation.
Super-resolution. We employ bicubic downsampling as the forward operator, as implemented in [48]. The learning rate is set to 12 for $\times$12 super-resolution and to 6 for $\times$8 super-resolution.
Motion Deblurring. A different motion blur kernel is created for each sample using the MotionBlur package [8], available via github, with kernel size 61 and intensity 0.5. The learning rate for our data term optimizer is set to $10^{-1}$.
Inpainting. For inpainting on FFHQ we always use the same rectangular mask at a fixed position, chosen such that it roughly masks out the right side of the face (Figure 7). For DIV2k we also use a fixed mask for all samples, consisting of six randomly generated rectangles (Figure 10).
Data. We use the publicly available Flickr Faces High Quality dataset [21], which is realeased under the Creative Commons BY 2.0 License and the DIV2K dataset [3], which is released under a research only license. For FFHQ we use the first 1000 samples of the evaluation dataset and for DIV2K we use the 800 training samples. We downscale both datasets to $768\times 768$ px by applying bicubic sampling so that the shorter edge of the frame has 768 px and apply central cropping afterwards.
A.5 Baselines
For comparability, all baselines use Stable Diffusion 3.5-Medium and the same task definitions as in A.4.
FlowDPS [23] The standard FlowDPS implementation [24] is applied with 50 NFE, a classifier-free guidance scale of 2, and step sizes of 15 for inpainting and 10 for all other tasks.
FlowChef [35] Additionally, [24] is employed for FlowChef as well, using 50 NFE, a classifier-free guidance scale of 2, and a step size of 1 for all tasks.
Repulsive Score Distillation (RSD) [54]. We implement RSD for flow-matching models by applying Proposition 1 with $\omega(t)=t$, resulting in a weighting term consistent with the original RSD approach. However, we omit the pixel-space augmentation as it negatively affected performance when combined with the SD3 VAE. Consistent with the original findings from RSD, we observed that incorporating the repulsive term improves sample diversity but reduces fidelity. Therefore, we set the repulsive term to 0 for all results presented in the table, employing it exclusively for comparing posterior variances.
ReSample [43] We re implement ReSample for flow-matching by setting $\bar{\alpha}_{t}=\frac{\left(1-t\right)^{2}}{t^{2}+\left(1-t\right)^{2}}$. Furthermore, we compute the hard-data consistency at every iteration as larger skip steps seem to harm performance. We set the learning rate of the data term optimizer to 15 for all inverse problems.
PSLD [38] Our attempt to adapt PSLD following [24]—using 500 NFE, a classifier-free guidance scale of 2, and step sizes of 1 ($\times$12 super-resolution), 0.5 ($\times$8 super-resolution and motion deblurring) and 0.1 (inpainting)—did not yield meaningful results.
A.6 Regularizer weighting
Figure 6 displays the mean and standard deviation of the conditional flow matching loss $\mathcal{L}_{CFM}$ as a function of $t$, estimated over 100 samples. The loss function starts with high values at $t=1$, decreases over time, but then starts to rise again, and when reaching $t\approx 0.2$ even exceeds its initial value . The rising loss when approaching $t=0$ is due, in part, to the increasing difficulty of distinguishing high-frequency image content from residual noise. Another factor is that near $t=0$ the model operates in a highly sensitive regime where small prediction errors can cause disproportionately large deviations from the target, making accurate flow estimation particularly challenging in the final stages of the trajectory. We therefore modulate the regularization term according to the model error. Different weighting functions for $f(\mathcal{L}_{CFM})$ could be chosen that fulfill the condition $\lambda_{\mathcal{R}(t=0)}=0$. We simply take the reciprocal of the model error $\lambda_{\mathcal{R}(t)}=\mathcal{L}_{CFM,t}^{-1}$ as the regularization weight while $t\geq 0.2$, then set it to 0 for $t<0.2$. An alternative would be to shift the reciprocal of $\mathcal{L}_{CFM,t}^{-1}$ by $\mathcal{L}_{CFM,t=0}^{-1}$, such that $\lambda_{\mathcal{R}(t)}=\mathcal{L}_{CFM,t}^{-1}-\mathcal{L}_{CFM,t=0}^{-1}$. In Table 3 we compare our default weighting with this variant, denoted as $\lambda_{shift}$.
Figure 6: The Flow-Matching loss over time $t$.
Table 3: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$. We compare different weighting functions $\lambda_{\mathcal{R}(t)}$ based on the model error
SR $\times$8 SR $\times$12 Motion Deblurring Inpainting
[…]
FFHQ 768$\times$768
$\lambda_{shift}$ $0.246$ $27.4$ $0.793$ $29.91$ $0.286$ $24.4$ $0.766$ $28.19$ $0.237$ $14.5$ $0.790$ $29.84$ $0.180$ $8.2$ $0.828$ $23.58$
Ours 0.213 235 303 759 574 90.21323530375957490.213\,235\,303\,759\,574\,90.213 235 303 759 574 9 13.266 211 509 704 5913.2662115097045913.266\,211\,509\,704\,5913.266 211 509 704 59 $0.777$ $29.54$ 0.270 564 854 145 050 050.270564854145050050.270\,564\,854\,145\,050\,050.270 564 854 145 050 05 16.201 229 095 458 98416.20122909545898416.201\,229\,095\,458\,98416.201 229 095 458 984 0.740 209 460 258 483 90.74020946025848390.740\,209\,460\,258\,483\,90.740 209 460 258 483 9 27.714 468 002 319 33627.71446800231933627.714\,468\,002\,319\,33627.714 468 002 319 336 0.236 155 018 210 411 070.236155018210411070.236\,155\,018\,210\,411\,070.236 155 018 210 411 07 10.732 836 723 327 63710.73283672332763710.732\,836\,723\,327\,63710.732 836 723 327 637 $0.772$ 29.611 394 882 202 1529.6113948822021529.611\,394\,882\,202\,1529.611 394 882 202 15 0.183 834 627 270 698 550.183834627270698550.183\,834\,627\,270\,698\,550.183 834 627 270 698 55 8.741 712 570 190 438.741712570190438.741\,712\,570\,190\,438.741 712 570 190 43 0.827 794 373 035 430 90.82779437303543090.827\,794\,373\,035\,430\,90.827 794 373 035 430 9 23.693 054 199 218 7523.6930541992187523.693\,054\,199\,218\,7523.693 054 199 218 75
DIV2K 768$\times$768
$\lambda_{shift}$ $0.379$ $30.4$ $0.625$ $23.58$ $0.434$ $37.5$ $0.522$ $21.40$ $0.337$ $25.8$ $0.664$ $24.54$ $0.151$ $9.0$ $0.819$ $23.79$
Ours 0.352 998 495 101 928 70.35299849510192870.352\,998\,495\,101\,928\,70.352 998 495 101 928 7 26.456 972 122 192 38326.45697212219238326.456\,972\,122\,192\,38326.456 972 122 192 383 $0.607$ $23.30$ 0.421 490 550 041 198 730.421490550041198730.421\,490\,550\,041\,198\,730.421 490 550 041 198 73 32.118 663 787 841 832.118663787841832.118\,663\,787\,841\,832.118 663 787 841 8 0.525 365 948 677 0630.5253659486770630.525\,365\,948\,677\,0630.525 365 948 677 063 21.385 475 158 691 40621.38547515869140621.385\,475\,158\,691\,40621.385 475 158 691 406 0.314 553 380 012 512 20.31455338001251220.314\,553\,380\,012\,512\,20.314 553 380 012 512 2 21.111 848 831 176 75821.11184883117675821.111\,848\,831\,176\,75821.111 848 831 176 758 0.652 780 890 464 782 70.65278089046478270.652\,780\,890\,464\,782\,70.652 780 890 464 782 7 24.444 496 154 785 15624.44449615478515624.444\,496\,154\,785\,15624.444 496 154 785 156 0.162 881 195 545 196 530.162881195545196530.162\,881\,195\,545\,196\,530.162 881 195 545 196 53 11.026 673 316 955 56611.02667331695556611.026\,673\,316\,955\,56611.026 673 316 955 566 0.814 652 204 513 549 80.81465220451354980.814\,652\,204\,513\,549\,80.814 652 204 513 549 8 23.752 101 898 193 3623.7521018981933623.752\,101\,898\,193\,3623.752 101 898 193 36
A.7 Effect of captioning
Given the diversity of DIV2k, we use DAPE [52] to generate captions for it and include them in the prompt A high quality photo of [DAPE caption]. For FFHQ we always prompt with A high quality photo of a face.. The effect of the text prompt is to increase the likelihood of our sample under the prior of the (pre-trained, frozen) image generator. For comparison, we also ran experiments without data specific captions, where we always used the generic prompt A high quality photo. Results are shown in Table 4
Table 4: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$. We compare our version with data-specific captions and a version without captions.
FFHQ DIV2K
[…]
wo captions 0.278 17.0 0.734 27.66 0.488 51.5 0.546 21.82
Ours 0.271 16.2 0.740 27.71 0.421 32.1 0.525 21.39
A.8 Additional Experimental Results
Table 5: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$ – Super-resolution (×$\times 8$ and ×$\times 12$).
SR ×$\times 8$ SR ×$\times 12$
[…]
FFHQ 768$\times$768
ReSample 0.400 $\pm$ 0.069 55.6 0.815 $\pm$ 0.051 26.37 $\pm$ 1.00 0.474 $\pm$ 0.078 80.3 0.786 $\pm$ 0.056 25.47 $\pm$ 1.16
FlowDPS 0.374 $\pm$ 0.107 40.3 0.756 $\pm$ 0.075 29.24 $\pm$ 2.04 0.413 $\pm$ 0.107 47.2 0.741 $\pm$ 0.074 28.05 $\pm$ 2.06
RSD 0.391 $\pm$ 0.079 51.7 0.776 $\pm$ 0.052 29.69 $\pm$ 2.04 0.462 $\pm$ 0.093 71.7 0.743 $\pm$ 0.059 28.11 $\pm$ 2.00
FlowChef 0.341 $\pm$ 0.083 30.5 0.760 $\pm$ 0.064 28.42 $\pm$ 2.22 0.373 $\pm$ 0.084 41.0 0.730 $\pm$ 0.068 27.00 $\pm$ 2.07
Ours 0.213 $\pm$ 0.056 13.3 0.777 $\pm$ 0.051 29.54 $\pm$ 2.02 0.271 $\pm$ 0.071 16.2 0.740 $\pm$ 0.058 27.71 $\pm$ 2.00
DIV2K 768$\times$768
ReSample 0.533 $\pm$ 0.130 55.0 0.625 $\pm$ 0.132 22.34 $\pm$ 2.27 0.643 $\pm$ 0.152 88.1 0.562 $\pm$ 0.151 20.85 $\pm$ 3.02
FlowDPS 0.476 $\pm$ 0.129 54.2 0.567 $\pm$ 0.139 23.01 $\pm$ 3.01 0.547 $\pm$ 0.139 67.4 0.528 $\pm$ 0.146 21.79 $\pm$ 2.94
RSD 0.539 $\pm$ 0.121 60.9 0.591 $\pm$ 0.124 23.45 $\pm$ 2.96 0.684 $\pm$ 0.137 95.7 0.523 $\pm$ 0.132 21.96 $\pm$ 2.86
FlowChef 0.490 $\pm$ 0.116 48.6 0.539 $\pm$ 0.137 21.84 $\pm$ 2.96 0.525 $\pm$ 0.118 55.3 0.492 $\pm$ 0.145 20.52 $\pm$ 2.85
Ours 0.353 $\pm$ 0.112 26.5 0.607 $\pm$ 0.127 23.30 $\pm$ 2.90 0.421 $\pm$ 0.131 32.1 0.525 $\pm$ 0.136 21.39 $\pm$ 2.67
Table 6: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$ – Motion deblurring and in-painting.
Motion Deblurring In-painting
[…]
FFHQ 768$\times$768
ReSample 0.457 $\pm$ 0.087 82.9 0.788 $\pm$ 0.058 25.45 $\pm$ 1.46 0.366 $\pm$ 0.053 70.8 0.827 $\pm$ 0.033 21.83 $\pm$ 1.68
FlowDPS 0.431 $\pm$ 0.117 49.5 0.732 $\pm$ 0.078 27.64 $\pm$ 2.20 0.344 $\pm$ 0.060 38.3 0.771 $\pm$ 0.048 19.19 $\pm$ 3.19
RSD 0.458 $\pm$ 0.098 77.3 0.743 $\pm$ 0.059 27.67 $\pm$ 2.47 0.478 $\pm$ 0.082 73.3 0.736 $\pm$ 0.048 21.97 $\pm$ 2.58
FlowChef 0.406 $\pm$ 0.093 35.0 0.716 $\pm$ 0.072 25.81 $\pm$ 2.61 0.394 $\pm$ 0.069 71.0 0.780 $\pm$ 0.051 18.18 $\pm$ 2.84
Ours 0.236 $\pm$ 0.070 10.7 0.772 $\pm$ 0.055 29.61 $\pm$ 2.24 0.184 $\pm$ 0.038 8.7 0.828 $\pm$ 0.029 23.69 $\pm$ 2.77
DIV2K 768$\times$768
ReSample 0.556 $\pm$ 0.146 79.7 0.617 $\pm$ 0.134 21.79 $\pm$ 2.52 0.285 $\pm$ 0.073 51.9 0.796 $\pm$ 0.067 22.68 $\pm$ 1.84
FlowDPS 0.558 $\pm$ 0.153 60.4 0.536 $\pm$ 0.148 21.88 $\pm$ 3.02 0.328 $\pm$ 0.103 34.2 0.692 $\pm$ 0.112 21.71 $\pm$ 2.67
RSD 0.638 $\pm$ 0.156 97.6 0.551 $\pm$ 0.136 22.10 $\pm$ 3.07 0.464 $\pm$ 0.112 63.9 0.678 $\pm$ 0.077 23.23 $\pm$ 2.21
FlowChef 0.561 $\pm$ 0.123 51.3 0.486 $\pm$ 0.148 19.90 $\pm$ 3.06 0.489 $\pm$ 0.148 86.3 0.659 $\pm$ 0.131 20.87 $\pm$ 2.65
Ours 0.315 $\pm$ 0.107 21.1 0.653 $\pm$ 0.121 24.44 $\pm$ 3.05 0.163 $\pm$ 0.053 11.0 0.815 $\pm$ 0.054 23.75 $\pm$ 2.74
We present the experimental results from the main paper in Table 5, now augmented with sample-wise standard deviations for all metrics except FID.
A.9 Additional Qualitative Examples
To illustrate the visual differences behind the error metrics, we present additional qualitative results for both FFHQ and DIV2k, comparing FLAIR with existing approaches. These examples complement the images in the main paper and highlight the visual fidelity, consistency, and robustness of our method across diverse scenes and different degradations. Figure 13 features a full sized version of the variance figure in section subsection 5.2.
Figure 7: Inpainting results on FFHQ. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces realistic, high-frequency details while previous works either fail to inpaint the region correctly or collapse to overly smooth solutions.
Figure 8: ×$\times 12$ super-resolution results on FFHQ. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
Figure 9: Motion de-blur results on FFHQ. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
Figure 10: Inpainting results on DIV2k. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces realistic, high-frequency details while previous works either fail to inpaint the region correctly or collapse to overly smooth solutions. Moreover they do not fit the data term (not inpainted region) very well.
Figure 11: ×$\times 12$ super-resolution results on DIV2k. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
Figure 12: Motion de-blur results results on DIV2k. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
Figure 13: Individual samples for x12 Super Resolution with zoom and std. FLAIR produces varied samples from the posterior. For superresoltion The variance is expected to be mostly in the high frequencies, because the data term limits low frequency variations. The baselines tend to predict very similar looking images with less detail.
A.10 Failure cases
We observe two main failure modes for FLAIR, see Figure 14. First, we find that super-resolution on DIV2k occasionally results in grainy textures, usually in regions with abundant high-frequency detail and complicated light transport. Potentially, this happens for images which do not have high probability under the prior. We do not observe those artifacts for the FFHQ dataset. Second, we observe a few instances where the strong generative prior hallucinates semantically inconsistent or misaligned structures – especially facial features.
Figure 14: Qualitative failure cases of FLAIR on DIV2k and FFHQ. Top row: grainy results from systematic error. Those errors potentially stem from a weak prior for those images. For example we do not observe them for the FFHQ dataset Bottom row: Semantically inconsistent failures. Sometimes the model lacks the ability to incorporate globally consistent semantics into its restorations.