### v2

# Solving Inverse Problems with FLAIR

Abstract
Flow-based latent generative models such as Stable Diffusion 3 are able to generate images with remarkable quality, even enabling photorealistic text-to-image generation. Their impressive performance suggests that these models should also constitute powerful priors for inverse imaging problems, but that approach has not yet led to comparable fidelity. There are several key obstacles: (i) the data likelihood term is usually intractable; (ii) learned generative models cannot be directly conditioned on the distorted observations, leading to conflicting objectives between data likelihood and prior; and (iii) the reconstructions can deviate from the observed data. We present FLAIR, a novel, training-free variational framework that leverages flow-based generative models as prior for inverse problems. To that end, we introduce a variational objective for flow matching that is agnostic to the type of degradation, and combine it with deterministic trajectory adjustments to guide the prior towards regions which are more likely under the posterior. To enforce exact consistency with the observed data, we decouple the optimization of the data fidelity and regularization terms. Moreover, we introduce a time-dependent calibration scheme in which the strength of the regularization is modulated according to off-line accuracy estimates. Results on standard imaging benchmarks demonstrate that FLAIR consistently outperforms existing diffusion- and flow-based methods in terms of reconstruction quality and sample diversity. Source code is available at https://inverseflair.github.io/.
1 Introduction
Flow-based generative models are at the core of modern image generators like Stable Diffusion or FLUX esser24icml . Beyond image generation based on text prompts, these models have emerged as powerful data-driven priors for a whole range of visual computing tasks. Their comprehensive representation of the visual world, learned from internet-scale training datasets, makes them an attractive alternative to traditional handcrafted image priors. Often, they can be used without any task-specific retraining.
While it is evident that a model capable of generating photorealistic images should be suitable as prior (a.k.a. regularizer) for inverse imaging problems, a practical implementation faces several challenges. On the one hand, flow-based models normally operate in the lower-dimensional latent space of a variational autoencoder (VAE), which means that the forward operator (the relationship between the observed, degraded image and the desired, clean target image) is no longer linear. On the other hand, the iterative nature of the generative process means that intermediate stages are corrupted with (time-dependent) random noise. Hence, one cannot explicitly evaluate their data likelihood, which renders the data term intractable. Moreover, learned generative models tend to overly favor regions of the training distribution that have a high sample density. For test samples that fall in low-density regions, the prior will have a too strong tendency to pull towards outputs with higher a-priori likelihood, compromising fidelity to the input observations.
Here, we propose flow-based latent adaptive inference with deterministic re-noising (FLAIR), a novel, training-free variational framework explicitly tailored to integrate flow-based latent diffusion models into inverse problem-solving. To the best of our knowledge, FLAIR is the first scheme that combines latent generative modeling, flow matching and variational inference into a unified formulation for inverse problems. Our main contributions are
  • A novel variational objective for inverse problems with flow-matching priors.
  • Deterministic trajectory adjustments guide the prior towards regions which are more consistent with the observed data.
  • Decoupled optimization of data and regularization terms, enabling hard data consistency.
  • A novel, time-dependent weighting scheme, calibrated via offline accuracy estimates, that adapts the regularization along the flow trajectory to match the changing reliability of the model’s predictions, ensuring robust inference.
Figure 1: Starting from the adjoint based initialization, we alternate between (i) regularizer updates via a flow-matching loss that aligns the velocity $u_{t}$ of the variational distribution with the learned velocity field $v_{\theta}$, and (ii) hard data consistency steps that project the current estimate onto the measurement manifold.
2 Related Work
Deep learning based priors. Deep learning–based methods typically follow one of two main approaches: they either directly learn an inverse mapping zb18 ; ja08 ; le17 ; becker2023neural ; Zhao_2023_CVPR , or aim to learn a suitable prior, either through non-generative approaches like unrolled optimization networks ham18 ; koef21 ; aga18 ; narnhofer2021bayesian or through generative models such as generative adversarial networks na19 ; bor17 ; shah2018solving , or diffusion ho2020denoising_ddpm ; song2020denoising_ddim ; song2021scorebased respectively flow based models lipman2023flow ; liu2022flow . The latter have demonstrated impressive performance in image generation tasks, sparking growing interest in leveraging them as priors for solving inverse problems, particularly through posterior sampling techniques.
Posterior sampling. Although incorporating the prior learned from a diffusion or flow-based model seems straightforward, problems arise due to the inherent time-dependent structure of diffusion models, which makes the likelihood term intractable chung2023diffusion . A variety of approaches have been proposed for diffusion-based posterior sampling zhang2025improving ; janati2025mixture ; moufad2024variational ; janati2024divide : enforcing the trajectory to stay on the respective noise manifold chung2023diffusion ; chung2022improving ; Zhao_2023_ICCV , applying an SVD to run diffusion in the spectral domain kawar2021snips , utilizing range-null space decomposition during the reverse diffusion process kawar2021snips , guidance by the pseudo-inverse of the forward operator song2023pseudoinverse .
Many prior methods perform well in pixel space but are difficult to apply in latent diffusion models due to VAE non-linearity or memory constraints. In order to circumvent this issue, the authors of ReSample song2024solving rely on enforcing hard data consistency through optimization and resampling during the reverse diffusion process. PSLD rout2023solving , introduces additional objectives terms to ensure that all gradient updates point to the same optima in the latent space. FlowChef patel2024flowchef incorporates guidance into the flow trajectory during inference, whereas FlowDPS kim2025flowdps separates the update step into two components: one for estimating the clean image and another for estimating the noise.
In contrast FLAIR follows another class of posterior sampling-based methods, which integrate diffusion priors with inverse problems by directly optimizing a variational objective that approximates the data posterior mardani2024a . This framework was recently extended by RSD zilberstein2025repulsive , which incorporates a repulsion mechanism to promote sample diversity and applied to latent diffusion models. A known issue with this type of optimization is mode collapse pooledreamfusion , which leads to blurry results for these methods. Our method targets this problem by introducing a deterministic trajectory adjustment.
3 Background
3.1 Inverse problems
In many imaging tasks, such as inpainting bertalmio2000image , super-resolution park2003super or tomographic reconstruction sidky2008image , one aims to recover a target signal $x\in\mathbb{R}^{n}$ from a distorted observation $y\in\mathbb{R}^{m}$. The observation is regarded as the result of applying a forward operator $\mathcal{A}:\mathbb{R}^{n}\mapsto\mathbb{R}^{m}$ to the target signal, corrupted by additive Gaussian noise $\nu\in\mathbb{R}^{m}$ with standard deviation $\sigma_{\nu}$.
$y=\mathcal{A}x+\nu.$ (1)
In most practical applications, the forward operator $\mathcal{A}$ is either non-invertible or severely ill-conditioned, making (1) generally ill-posed.
Variational methods solve ill-posed inverse problems by minimizing an energy functional
$\mathcal{E}(x,y)=\mathcal{D}(x,y)+\mathcal{R}(x).$ (2)
to recover the solution.
Interpreted probabilistically via Bayes’ theorem, the posterior distribution $p(x|y)$ is proportional to the product $p(y|x)p(x)$. In the negative log-domain, this yields the data term $\mathcal{D}(x,y)=-\log p(y|x)$ and the regularizer $\mathcal{R}(x)=-\log p(x)$. Handcrafted priors based on regularity assumptions like sparsity rof92 ; sa01 ; chpo11 ; da04 ; ho20_ip_review have long been the standard, but have largely been replaced by deep learning-based methods in modern data-driven schemes.
3.2 Flow based priors
Models based on flow matching lipman2023flow learn a time-dependent vector field $v_{\theta}(x_{t},t)$ that continuously transforms samples from a simple initial distribution $p_{1}(x)$ to a complex target data distribution $p_{0}(x)$. Formally, this transformation is described by solving the ordinary differential equation (ODE):
$\frac{\operatorname{d}\!{}}{\operatorname{d}\!{t}}\psi_{t}(x)=v_{\theta,t}(\psi_{t}(x)),\quad t\in[0,1],$ (3)
where $\psi_{t}(x)$ represents the trajectory of a sample, evolving smoothly from an initial value drawn at $t=1$ toward a target value at $t=0$.
Since the integrated ODE path maps the simple distribution $p_{1}(x)$ to the complex target $p_{0}(x)$, the learned flow-based model captures the structure of the data and can therefore serve as a powerful prior for solving inverse problems. To make this approach tractable for high-resolution data, we adopt the latent diffusion model (LDM) framework rombach2022high , which shifts the generative process to a lower-dimensional latent space using a pretrained autoencoder with encoder $E:\mathbb{R}^{n}\mapsto\mathbb{R}^{d}$ and decoder $D:\mathbb{R}^{d}\mapsto\mathbb{R}^{n}$, where $d\ll n$. However, applying such priors to inverse problems introduces challenges, as the non-linearity of the VAE disrupts the linear relationship between measurements and the target signal, resulting in a nonlinear forward operator.
3.3 Variational flow sampling
To solve inverse problems from a Bayesian perspective, we aim to sample from the posterior
$p(x_{0}|y)\propto p(y|x_{0})p(x_{0}),$ (4)
where the likelihood is given by $p(y|x_{0})=\mathcal{N}(\mathcal{A}x_{0},\sigma^{2}\mathrm{Id})$, and $p(x_{0})$ represents the prior modeled by the flow-based generative model.
Inspired by previous work mardani2024a ; zilberstein2025repulsive we introduce a variational distribution $q(x_{0}|y)=\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ to approximate the true posterior $p(x_{0}|y)$, by minimizing their Kullback–Leibler divergence:
$q(x_{0}|y)\in\arg\min_{q(x_{0}|y)}\text{KL}(q(x_{0}|y)\|p(x_{0}|y)).$ (5)
Rewriting the KL divergence by means of the variational lower bound leads to:
$KL(q(x_{0}|y)\|p(x_{0}|y))=-\underbrace{\mathbb{E}_{q(x_{0}|y)}[\log p(y|x_{0})]}_{\text{$\mathcal{D}(x,y)$}}+\underbrace{KL(q(x_{0}|y)\|p(x_{0}))}_{\text{$\mathcal{R}(x)$}}+\underbrace{\log p(y)}_{\text{const}}.$ (6)
Since a single Gaussian cannot capture a multi-modal posterior, we simplify to a deterministic approximation, setting $\sigma_{x}^{2}=0$. Equivalently, this corresponds to a single-particle approximation in the sense of Stein variational methods liu2016stein . As shown in song2021maximum , rewriting Equation 6 under this approximation and extending it to the time-dependent noisy posterior yields:
$\arg\min_{q(x_{0}|y)}\underbrace{\mathbb{E}_{q(x_{0}|y)}\left[\frac{\|y-f(\mu_{x})\|^{2}}{2\sigma_{\nu}^{2}}\right]}_{\text{$\mathcal{D}(x,y)$}}+\underbrace{\int_{0}^{T}\omega(t)~\mathbb{E}_{q(x_{t}|y)}\left[\left\|\nabla_{x}\log q(x_{t}|y)-\nabla_{x}\log p(x_{t})\right\|^{2}\right]dt}_{\text{$\mathcal{R}(x)$}}$ (7)
The first term in Equation 7 describes the data term $\mathcal{D}(x,y)$ and the second the regularizer $\mathcal{R}(x)$, where the integral ensures optimization over the entire diffusion trajectory. Notably, the latter constitutes a weighted score-matching objective, where $\nabla_{x}\log p(x_{t})$ represents the score function song2021scorebased , which may be extracted from a pretrained diffusion or flow model.
The score of the noisy variational distribution depends on the forward diffusion process and can be computed analytically.
Note that for $\omega(t)=\beta(t)/2$ the weighted score-matching loss recovers the gradient of the diffusion model’s evidence lower bound, so that optimizing it yields the maximum likelihood estimate of the data distribution song2021maximum . However, optimizing Equation 7 is costly, as it requires computing the gradient through the flow model. As shown in wang2023prolificdreamer this can be circumvented by reformulating the regularizer in terms of the Wasserstein gradient flow:
[…]
Note that optimizing only the regularization term, without the data term, at test time is equivalent to the objective of Score Distillation Sampling (SDS) pooledreamfusion .
4 Method
Flow Formulation. The variational formulation in Equation 7 is formulated for the score, but can be reformulated into a denoising or $\epsilon_{\theta}$ parameterization song2021scorebased ; mardani2024a . However, we are interested in a variational objective that depends on the velocity field $v_{\theta}(x_{t},t)$, which characterizes the probabilistic trajectory that connects the noise and data distributions.
Proposition 1.
We propose to replace the score-based regularizer in the standard variational objective with a flow matching formulation, resulting in the following objective function:
$\arg\min_{q(x_{0}|y)}\underbrace{\mathbb{E}_{q(x_{0}|y)}\left[\frac{\|y-f(\mu_{x})\|^{2}}{2\sigma_{\nu}^{2}}\right]}_{\text{$\mathcal{D}(x,y)$}}+\underbrace{\int_{0}^{T}\lambda_{\mathcal{R}}(t)~\mathbb{E}_{q(x_{t}|y)}\left[\left\|v_{\theta}(x_{t},t)-u_{t}(x_{t}|\epsilon)\right\|^{2}\right]dt}_{\text{$\mathcal{R}(x)$}}$ (9)
$\nabla_{\mu_{x}}\mathcal{R}(x)=\mathbb{E}_{t,q(x_{t}|y)}\left[\lambda_{\mathcal{R}}(t)v_{\theta}(x_{t},t)-u_{t}(x_{t}\mid\epsilon)\right]$ (10)
The flow-matching term that defines the regularizer arises by reparameterizing the variational distribution to $q(x_{t}|y)=\mathcal{N}((1-t)\mu_{x},t^{2}\,I)$. This corresponds to sampling via the deterministic map $\psi_{t}(x_{0}\mid\epsilon)=(1-t)\,x_{0}+t\,\epsilon$, with $\epsilon\sim\mathcal{N}(0,I)$. By reformulating the score in terms of the target velocity field $u_{t}$, we get:
$\nabla_{x}\log q(x_{t}|y)=-\frac{(1-t)u_{t}(x_{t}|\epsilon)+x_{t}}{t}$ (11)
For the learned velocity $v_{\theta}(x_{t},t)$ a similar approximation holds – for a full derivation, see the supplementary material subsection A.3.
$v_{\theta}(x_{t},t)\approx\frac{-t\nabla_{x}\log p(x_{t})-x_{t}}{1-t}$ (12)
We can therefore approximate the score of the noisy prior with our learned velocity field $v_{\theta}$
$\nabla_{x}\log p(x_{t})\approx-\frac{(1-t)v_{\theta}(x_{t},t)+x_{t}}{t}.$ (13)
Hard Data Consistency. Existing variational posterior sampling approaches mardani2024a ; zilberstein2025repulsive impose soft constraints on the data fidelity term $\mathcal{D}(x,y)$. In contrast, recent work song2024solving has demonstrated that, when sampling from latent diffusion models, enforcing hard data consistency generally leads to better reconstructions with improved visual fidelity. Our method shares this motivation, but differs in that we optimize over a variational distribution, i.e., we compute $\min~\mathbb{E}_{q(x_{0}|y)}[-\log p(y|x_{0})]$. An additional advantage of this variational setup is that it allows us to initialize the optimization variable with an adjoint based initialization $\mu_{x}=E(A^{\top}y)$, with $E$ being the encoder of the VAE and $A^{\top}$ the adjoint of the linear forward operator in pixel space. Other initialization strategies are also possible.
Accuracy Calibration. As our framework evaluates the trajectory at each time step, we aim to weight the regularizer’s contribution according to its reliability. The difficulty of the prediction task has been shown to depend on the network parameterization, as well as on the specific time step $t$ Karras2022edm . Since the regularization term $\mathcal{R}(x)$ in our approach is equivalent to the training objective of the pre-trained flow model, we can easily weight it by the expected model error, which we calibrate on a small set of images. Specifically, we sample $N$ calibration images and compute the conditional flow matching objective for 100 linearly spaced time steps between 0 and 1, then average the error over all images to obtain the expected model error at each time step. Different functions of the model error can be chosen as weight for the regularizer. We choose:
$\lambda_{\mathcal{R}}(t)=\frac{1}{N}\left(\sum_{i=1}^{N}\left\|v_{\theta}(x^{(i)}_{t},t)-u_{t}(x^{(i)}_{t}\mid\epsilon)\right\|^{2}\right)^{-1}$ (14)
and set $\lambda_{\mathcal{R}}(t)=0$ for all $t<0.2$, since the accuracy of SD3 is heavily degraded for low noise levels.
Deterministic Trajectory Adjustment. Score distillation sampling relies on the assumption that $x_{t}=(1-t)\mu_{x}+t\epsilon$ lies in a region of the learned prior that has reasonably high support/density. In practice, this is not always the case. When not tightly conditioned (usually with extensive text prompts), even the best available diffusion models assign low density to many plausible regions of the latent space, leading to bad gradient steps. Therefore, we increase the probability of $p(x_{t})$ by additionally conditioning $x_{t}$ on the estimated "end-point" $\mu$.
Proposition 2.
We introduce a reparameterized variational distribution with a mean that linearly interpolates between the posterior mean $\mu_{x}$ and a model-guided  $\hat{x}_{1}$:
$q(x_{t}\mid y)=\mathcal{N}\left((1-t)\mu_{x}+t\alpha\hat{x}_{1},\;t^{2}(1-\alpha^{2})I\right),$ (15)
where $\hat{x}_{1}=x_{t+\delta t}+(1-t-\delta t)v_{\theta}(x_{t+\delta t},t+\delta t)$ is a single-step velocity-based predictor, and $\alpha\in[0,1]$ controls the trade-off between deterministic guidance and random noise. This reparameterization induces the following reference velocity field:
$u_{t}(x_{t}\mid\epsilon)=\frac{\alpha\hat{x}_{1}+\sqrt{1-\alpha^{2}}\epsilon-x_{t}}{1-t}.$ (16)
Intuitively, changing the formulation in this manner ensures that the model relocates the sample to its expected position on the learned manifold rather than injecting arbitrary noise, which could drive it in a direction that has high prior likelihood but is not consistent with the observation. To further encourage exploration and avoid collapsing onto the trajectory of the adjoint measurement, we inject an additional stochastic component $\epsilon$ during this process. A full derivation can be found in the supplementary material, subsection A.3.
4.1 Algorithm
The following pseudo-code summarizes our method, integrating all the components discussed above.
We adapt the standard scheme mardani2024a ; zilberstein2025repulsive of linearly traversing time in a descending manner and stop at $t=0.2$ as explained in section 4. We choose $\alpha=1-t$. Gradient updates to enforce hard data consistency are performed using stochastic gradient descent. For further implementation details and ablations, see subsection A.4
Input: $\mu_{x}=\mu_{init}$, $\lambda_{R}$, $\alpha$, $y$, $\mathcal{A}$, $v_{\theta}$
Output: $\mu_{x}$
$\hat{\epsilon}\sim\mathcal{N}(0,I)$;
$\triangleright$ initial noise sample
for $t\leftarrow 1$ to 0 by $\Delta t$ do
 $x_{t}\leftarrow(1-t)\,\mu_{x}+t\,\hat{\epsilon}$;
  $\triangleright$ sample noisy latent
  $u_{t}(x_{t}\mid\hat{\epsilon})\leftarrow\dfrac{\hat{\epsilon}-x_{t}}{1-t}$;
[…]
 $\mu_{x}\leftarrow\mu_{x}-\lambda_{R}\,\nabla_{\mu_{x}}R$;
  $\triangleright$ update w.r.t. regularizer
 $\mu_{x}\leftarrow\arg\min_{\mu_{x}}\|y-\mathcal{A}(\mu_{x})\|^{2}$;
  $\triangleright$ hard data consistency
 $\epsilon\sim\mathcal{N}(0,I)$;
 $\hat{x}_{1}\leftarrow x_{t}+(1-t)\,v_{\theta}(x_{t},t)$;
  $\triangleright$ predict deterministic noise
  $\hat{\epsilon}\leftarrow\alpha\,\hat{x}_{1}+\sqrt{1-\alpha^{2}}\,\epsilon$;
  $\triangleright$ update noise estimate
Algorithm 1 The FLAIR solver for inverse imaging problems
5 Experiments
We evaluate the performance of FLAIR in a variety of inverse imaging tasks and compare it against several baselines, using the SD3 backbone without any fine-tuning. We used several metrics including SSIM wang2004image , LPIPS zhang2018unreasonable and patchwise FID heusel2017gans (pFID) to comprehensively assess the perceptual and quantitative quality of the reconstructions. FID is computed using InceptionV3 features on patches of 256x256 resolution. All experiments were performed on a NVidia RTX 4090 GPU with 24GB of VRAM. For completeness we also show PSNR values, but point out that the metric is not well suited for our purposes: PSNR favors the posterior mean, while the goal of the variational approach is to sample from the posterior distribution. Accordingly, PSNR is known to prefer over-smoothed, blurry outputs over sharp ones blau2018perception . To demonstrate that our model can also produce accurate MMSE estimates, we performed ensemble predictions by running posterior sampling eight times and averaging the results. As shown in subsection A.11, ensembling improves PSNR values while reducing LPIPS. This confirms that our samples are distributed around the posterior mean. Moreover, it shows that results closer to the posterior mean – such as those produced by baseline methods – are perceptually farther from the ground truth (in LPIPS) compared to our samples.
5.1 Setup
Datasets. We utilize two high-resolution image datasets: FFHQ karras2019style and DIV2K Agustsson_2017_CVPR_Workshops . FFHQ consists of 70k diverse face images at 1024×1024 resolution of which we take the first 1000 samples. It is covering variations in age, pose, lighting, and ethnicity. DIV2K contains 800 high-quality images in 2K resolution that span a range of natural scenes with varied textures and structures.
Baselines. Our method is benchmarked against several recent inverse imaging solvers based on posterior sampling. Specifically, we compare to ReSample song2024solving , FlowDPS kim2025flowdps , FlowChef patel2024flowchef , and RSD zilberstein2025repulsive . The latter is used without repulsive term as it delivers better results. To ensure a fair and meaningful comparison, all methods are evaluated with the same number of function evaluations.
Problem Setting. We run and evaluate all methods at a fixed output resolution of $\times768$ pixels. For single image super-resolution, we consider scaling factors of $8\times$ and $12\times$. The corresponding low-resolution inputs are generated by bicubic downsampling. Motion blur is simulated with a blur kernel of size 61. For box inpainting, we mask large, continuous rectangles that cover approximately one third of the observation. All synthesized observations are corrupted with additive Gaussian noise, with standard deviation $\sigma_{\nu}$ of 0.5%.
For inference on the FFHQ dataset, we use a predefined text prompt of the form "A high quality photo of a face", and for DIV2k "A high quality photo of" concatenated with an image-specific description retrieved by applying DAPE Wu_2024_CVPR to the observation.
5.2 Experimental Results
Inverse Problems. Our experiments clearly demonstrate that FLAIR outperforms existing flow-based approaches in terms of all perceptual metrics, see Table 1.
In the case of image inpainting, our method produces high-quality reconstructions that fully leverage the power of the generative model and blend naturally into the surrounding context, avoiding degradations and artifacts that we observe in the baselines. In particular, FlowDPS tends to produce implausible textures in the inpainted regions, while FlowChef regularly fails to generate semantically consistent content at all.
For single-image super-resolution, FLAIR consistently delivers the most perceptually convincing and realistic outputs. Notably, the FID scores remain low for both $\times 8$ and $\times 12$ magnification, indicating an effective usage of the generative prior to overcome the increasing ill-posedness. Again, FlowDPS suffers from blur and low texture quality, whereas FlowChef tends to lose semantic coherence.
In motion deblurring, FLAIR also restores sharper and semantically more credible content than competing approaches, which often suffer from residual blur or inconsistent details. The boost in reconstruction quality is quantitatively reflected by all metrics, confirming that FLAIR reconstructs images with high fidelity. For further qualitative examples, see subsection A.13.
Refer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurring
Figure 2: Qualitative comparison. FLAIR produces posterior samples of high perceptual quality while maintaining high data likelihood. Best viewed zoomed in.
Table 1: Quantitative results with 50 NFE and $\sigma_{\nu}=0.5\%$.
SR $\times$8 SR $\times$12 Motion Deblurring Inpainting
[…]
FFHQ $\times768$
ReSample 0.400 331 109 762 191 80.400\,331\,109\,762\,191\,8 55.586 227 416 992 1955.586\,227\,416\,992\,19 0.814 689 397 811 889 60.814\,689\,397\,811\,889\,6 26.367 465 972 900 3926.367\,465\,972\,900\,39 0.473 518 252 372 741 70.473\,518\,252\,372\,741\,7 80.268 859 863 281 2580.268\,859\,863\,281\,25 0.785 562 992 095 947 30.785\,562\,992\,095\,947\,3 25.469 425 201 416 01625.469\,425\,201\,416\,016 0.457 411 199 808 120 70.457\,411\,199\,808\,120\,7 82.892 143 249 511 7282.892\,143\,249\,511\,72 0.787 788 689 136 505 10.787\,788\,689\,136\,505\,1 25.446 083 068 847 65625.446\,083\,068\,847\,656 0.365 910 947 322 845 460.365\,910\,947\,322\,845\,46 70.832 069 396 972 6670.832\,069\,396\,972\,66 0.827 21.827 827 453 613 2821.827\,827\,453\,613\,28
FlowDPS 0.374 372 541 904 449 460.374\,372\,541\,904\,449\,46 $38.528$ 0.756 326 615 810 394 30.756\,326\,615\,810\,394\,3 29.238 021 850 585 93829.238\,021\,850\,585\,938 0.413 043 200 969 696 040.413\,043\,200\,969\,696\,04 44.0 0.741 228 818 893 432 60.741\,228\,818\,893\,432\,6 28.05 0.431 281 954 050 064 10.431\,281\,954\,050\,064\,1 $54.300$ 0.731 617 748 737 335 20.731\,617\,748\,737\,335\,2 27.637 281 417 846 6827.637\,281\,417\,846\,68 0.344 42.5 0.770 720 958 709 716 80.770\,720\,958\,709\,716\,8 19.191 648 483 276 36719.191\,648\,483\,276\,367
RSD 0.391 473 561 525 344 850.391\,473\,561\,525\,344\,85 51.667 427 062 988 2851.667\,427\,062\,988\,28 0.776 033 043 861 389 20.776\,033\,043\,861\,389\,2 29.690 019 607 543 94529.690\,019\,607\,543\,945 0.462 387 174 367 904 660.462\,387\,174\,367\,904\,66 71.734 550 476 074 2271.734\,550\,476\,074\,22 0.743 28.108 963 012 695 31228.108\,963\,012\,695\,312 0.457 823 425 531 387 330.457\,823\,425\,531\,387\,33 77.319 137 573 242 1977.319\,137\,573\,242\,19 0.743 024 885 654 449 50.743\,024\,885\,654\,449\,5 27.67 0.478 281 795 978 546 140.478\,281\,795\,978\,546\,14 73.293 403 625 488 2873.293\,403\,625\,488\,28 0.735 569 655 895 233 20.735\,569\,655\,895\,233\,2 21.97
FlowChef 0.341 30.5 0.760 072 886 943 817 10.760\,072\,886\,943\,817\,1 28.424 997 329 711 91428.424\,997\,329\,711\,914 0.373 $46.5$ 0.730 421 423 912 048 30.730\,421\,423\,912\,048\,3 26.996 358 871 459 9626.996\,358\,871\,459\,96 0.406 40.2 0.716 119 468 212 127 70.716\,119\,468\,212\,127\,7 25.809 129 714 965 8225.809\,129\,714\,965\,82 0.394 163 548 946 380 60.394\,163\,548\,946\,380\,6 $69.837$ 0.780 103 147 029 876 70.780\,103\,147\,029\,876\,7 18.184 221 267 700 19518.184\,221\,267\,700\,195
Ours 0.213 235 303 759 574 90.213\,235\,303\,759\,574\,9 13.266 211 509 704 5913.266\,211\,509\,704\,59 0.777 29.54 0.270 564 854 145 050 050.270\,564\,854\,145\,050\,05 16.201 229 095 458 98416.201\,229\,095\,458\,984 0.740 209 460 258 483 90.740\,209\,460\,258\,483\,9 27.714 468 002 319 33627.714\,468\,002\,319\,336 0.236 155 018 210 411 070.236\,155\,018\,210\,411\,07 10.732 836 723 327 63710.732\,836\,723\,327\,637 0.772 29.611 394 882 202 1529.611\,394\,882\,202\,15 0.183 834 627 270 698 550.183\,834\,627\,270\,698\,55 8.741 712 570 190 438.741\,712\,570\,190\,43 0.827 794 373 035 430 90.827\,794\,373\,035\,430\,9 23.693 054 199 218 7523.693\,054\,199\,218\,75
DIV2K $\times768$
ReSample $0.533$ 54.997 371 673 583 9854.997\,371\,673\,583\,98 $0.625$ $22.34$ 0.642 829 596 996 307 40.642\,829\,596\,996\,307\,4 88.051 658 630 371 188.051\,658\,630\,371\,1 0.561 500 668 525 695 80.561\,500\,668\,525\,695\,8 $20.85$ 0.555 853 188 037 872 30.555\,853\,188\,037\,872\,3 79.654 029 846 191 479.654\,029\,846\,191\,4 0.617 21.792 427 062 988 2821.792\,427\,062\,988\,28 0.285 51.859 077 453 613 2851.859\,077\,453\,613\,28 0.796 22.676 401 138 305 66422.676\,401\,138\,305\,664
FlowDPS 0.476 $44.414$ 0.567 384 541 034 698 50.567\,384\,541\,034\,698\,5 23.005 729 675 292 9723.005\,729\,675\,292\,97 0.547 238 528 728 485 10.547\,238\,528\,728\,485\,1 $54.046$ 0.528 21.79 0.558 460 950 851 440 40.558\,460\,950\,851\,440\,4 $65.460$ 0.535 993 039 608 001 70.535\,993\,039\,608\,001\,7 21.877 470 016 479 49221.877\,470\,016\,479\,492 0.327 635 318 040 847 80.327\,635\,318\,040\,847\,8 29.2 0.691 909 790 039 062 50.691\,909\,790\,039\,062\,5 21.714 540 481 567 38321.714\,540\,481\,567\,383
RSD 0.538 948 655 128 4790.538\,948\,655\,128\,479 60.930 324 554 443 3660.930\,324\,554\,443\,36 0.591 312 229 633 331 30.591\,312\,229\,633\,331\,3 23.445 989 608 764 6523.445\,989\,608\,764\,65 0.683 576 822 280 8830.683\,576\,822\,280\,883 95.726 249 694 824 2295.726\,249\,694\,824\,22 0.523 275 852 203 369 10.523\,275\,852\,203\,369\,1 21.964 975 357 055 66421.964\,975\,357\,055\,664 0.637 668 132 781 982 40.637\,668\,132\,781\,982\,4 97.646 888 732 910 1697.646\,888\,732\,910\,16 0.551 110 327 243 804 90.551\,110\,327\,243\,804\,9 22.10 0.463 778 406 381 607 060.463\,778\,406\,381\,607\,06 63.870 880 126 953 12563.870\,880\,126\,953\,125 0.677 523 434 162 139 90.677\,523\,434\,162\,139\,9 23.23
FlowChef 0.490 322 709 083 557 130.490\,322\,709\,083\,557\,13 36.5 0.539 492 428 302 764 90.539\,492\,428\,302\,764\,9 21.842 645 645 141 621.842\,645\,645\,141\,6 0.525 43.8 0.492 154 121 398 925 80.492\,154\,121\,398\,925\,8 20.521 999 359 130 8620.521\,999\,359\,130\,86 0.561 49.6 0.486 047 297 716 140 750.486\,047\,297\,716\,140\,75 19.899 375 915 527 34419.899\,375\,915\,527\,344 0.488 857 418 298 721 30.488\,857\,418\,298\,721\,3 $58.339$ 0.659 262 299 537 658 70.659\,262\,299\,537\,658\,7 20.874 286 651 611 32820.874\,286\,651\,611\,328
Ours 0.352 998 495 101 928 70.352\,998\,495\,101\,928\,7 26.456 972 122 192 38326.456\,972\,122\,192\,383 0.607 23.30 0.421 490 550 041 198 730.421\,490\,550\,041\,198\,73 32.118 663 787 841 832.118\,663\,787\,841\,8 0.525 365 948 677 0630.525\,365\,948\,677\,063 21.385 475 158 691 40621.385\,475\,158\,691\,406 0.314 553 380 012 512 20.314\,553\,380\,012\,512\,2 21.111 848 831 176 75821.111\,848\,831\,176\,758 0.652 780 890 464 782 70.652\,780\,890\,464\,782\,7 24.444 496 154 785 15624.444\,496\,154\,785\,156 0.162 881 195 545 196 530.162\,881\,195\,545\,196\,53 11.026 673 316 955 56611.026\,673\,316\,955\,566 0.814 652 204 513 549 80.814\,652\,204\,513\,549\,8 23.752 101 898 193 3623.752\,101\,898\,193\,36
Posterior Variance. To demonstrate that FLAIR does not suffer from mode collapse, we assess the posterior variance $\operatorname*{Var}[x|y]$ for the task of $\times 12$ Super Resolution, by drawing 32 samples for a fixed observation $y$ and computing their pixel-wise variance. We conduct that experiment for our FLAIR approach, for RSD with repulsive term, and for FlowDPS kim2025flowdps . The example in Figure 3 illustrates that FLAIR has the highest sample diversity, which is also reflected in the corresponding variance maps. Notably, the sample variance is concentrated in regions with high-frequency textures. This indicates that our method reliably reconstructs the posterior, whose low-frequency part is, in the super-resolution setting, tightly constrained by the likelihood term.
Figure 3: Zoomed-in reconstructions for x12 Super Resolution. We show posterior samples (col. 1–4) of FLAIR, FlowDPS, and RSD, posterior mean and standard deviation (over 32 samples, col. 5,6). 0 Refer to caption $0.16$
Editing. Beyond image restoration, we observe that our method also performs remarkably well for text-based image editing, simply presenting suitable target prompts during inpainting. Figure 4 illustrates a variety of edited images generated from the same photograph with the help of the depicted masks and prompts.
Refer to captionOriginal image Refer to caption"man wearing aviator glasses."Refer to caption"man wearing a wizard hat"Refer to caption"man with a Mike Tyson facial tattoo"Refer to caption"clown with a red nose,red makeup and a ruff"
Figure 4: Edited images shown alongside original, with prompts: "A high resolution portrait of a…"
[…]
Figure 5: Qualitative comparison. FLAIR in pixel space produces posterior samples of high perceptual quality while maintaining high data likelihood as well. Best viewed zoomed in.
Table 2: Quantitative results with 50 NFE and $\sigma_{\nu}=0.5\%$ – In-painting and Super-resolution ($\times 8$).
Inpainting SR $\times 8$
[…]
DDNM 0.158 $\pm$ 0.042 26.9 0.732 $\pm$ 0.037 18.31 $\pm$ 2.94 0.199 $\pm$ 0.052 31.9 0.635 $\pm$ 0.079 23.59 $\pm$ 1.64
DPS 0.195 $\pm$ 0.064 30.2 0.689 $\pm$ 0.077 20.49 $\pm$ 2.81 0.172 $\pm$ 0.058 27.8 0.658 $\pm$ 0.088 24.59 $\pm$ 2.04
$M$ 0.161 $\pm$ 0.054 28.8 0.728 $\pm$ 0.062 20.59 $\pm$ 3.37 0.172 $\pm$ 0.051 29.1 0.669 $\pm$ 0.083 24.65 $\pm$ 1.97
[…]
FLAIR 0.097 $\pm$ 0.035 14.2 0.831 $\pm$ 0.031 21.87 $\pm$ 2.66 0.143 $\pm$ 0.039 22.9 0.712 $\pm$ 0.076 25.93 $\pm$ 1.96
5.3 Ablation Studies
We systematically analyze the impact of key design choices in our method. Specifically, we ablate the deterministic trajectory adjustment, the use of hard data consistency, and the calibration of the regularizer weight for $\times 12$ super-resolution, using a subset of 100 samples from the FFHQ and DIV2K datasets. Quantitative and qualitative results are shown in Table 3 and Figure 6, respectively.
Hard Data Consistency (HDC). Dropping the hard data consistency degrades both metrics, with PSNR being particularly affected due to poorer alignment with the input observation, which is also evident in the visual example: the reconstruction is plausible but deviates from the observation.
Deterministic Trajectory Adjustment (DTA). The biggest performance drop compared to the full setup occurs when removing the deterministic trajectory adjustment, as random noise sampling harms the gradient updates in low-density regions of the prior. The reconstruction appears overly smooth and lacks texture details.
Calibrated Regularizer Weight (CRW). Replacing our calibrated regularizer weight with $\lambda_{\mathcal{R}}(t)=t$ also has a strong impact on perceptual quality: the result is visibly blurred if one ignores the changing accuracy of the regularizer along the flow trajectory.
Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗
Figure 6: Qualitative samples from the ablation study on $\times$12 Super Resolution.
Table 3: Ablation study for $\times$12 super-resolution on DIV2K and FFHQ. Model components are individually switched on or off.
HDC DTA CRW FFHQ DIV2K
LPIPS ↓ PSNR ↑ LPIPS ↓ PSNR ↑
0.259 27.45 0.427 21.05
0.297 27.17 0.467 20.82
0.432 27.20 0.622 21.69
0.363 28.58 0.583 21.98
0.392 28.33 0.605 21.99
Legend. HDC: Hard Data Consistency; DTA: Deterministic Trajectory Adjustment;
CRW: Calibrated Regularizer Weight. ✓ = included, ✗ = ablated.
6 Conclusion and Limitations
We have presented FLAIR, a training-free variational framework for inverse problems that uses a flow-based generative model as its image prior. By combining the power of (latent) flow-based models with a principled reconstruction of the posterior distribution, FLAIR addresses key limitations of existing methods. First, it is able to target the generation towards images, which match the observation, by aiding the degradation-agnostic flow matching loss with deterministic noise vectors. Second, it enables hard data consistency without sacrificing sample diversity, by decoupling the data consistency constraint from the regularization, while adaptively reweighting the latter according to its expected accuracy, calibrated offline. Experiments with different image datasets and tasks confirm that FLAIR consistently achieves higher reconstruction quality than existing baselines based on either flow matching or denoising diffusion. Notably, our proposed method achieves, at the same time, excellent perceptual quality, close adherence to the input observations, and high sample diversity.
Evidently, FLAIR inherits the limitations of the underlying generative model. These include biases caused by the selection of training data, constraints w.r.t. the output resolution, and a limited ability to recover out-of-distribution modes. Furthermore, our approach introduces additional hyper-parameters needed to control the deterministic trajectory adjustment. We note that high fidelity image restoration methods can potentially be misused for unethical image manipulations.
Acknowledgments This work was funded, in part, by Huawei Technologies Oy (Finland) Co. Ltd.
Supplementary Material
In the following, we provide detailed line-by-line derivations of the mathematical formulations used in the paper, as well as additional implementation details and experimental results.
Appendix A Derivations
A.1 Derivation of flow-based variational formulation
The linear conditional flow and it’s corresponding velocity are defined as:
$\psi_{t}(x_{0}\mid\epsilon)=(1-t)\,x_{0}+t\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)\;,$ (17)
$u_{t}(x_{t}|\epsilon)=\frac{\operatorname{d}\!{\psi_{t}}}{\operatorname{d}\!{t}}(\psi_{t}^{-1}(x_{t}|\epsilon)|\epsilon)\;.$ (18)
The score of the noisy variational distribution can be analytically computed with:
$q(x_{t}|y)=\mathcal{N}((1-t)\mu_{x},t^{2}\,I)\;,$ (19)
$\nabla_{x_{t}}\log q(x_{t}|y)=-\frac{\epsilon}{t}\;.$ (20)
We compute $\frac{\operatorname{d}\!{\psi_{t}}}{\operatorname{d}\!{t}}(x_{0}|\epsilon)=-x_{0}+\epsilon$ and $\psi_{t}^{-1}(x_{t}|\epsilon)$ and insert it into Equation 18:
$u_{t}(x_{t}|\epsilon)=\frac{\epsilon-x_{t}}{1-t}\;.$ (21)
Solving Equation 21 for $\epsilon$ and inserting in Equation 20 gives:
$\nabla_{x_{t}}\log q(x_{t}|y)=-\frac{(1-t)u_{t}(x_{t}|\epsilon)+x_{t}}{t}\;.$ (22)
For the learned velocity $v_{\theta}(x_{t},t)$ a similar approximation holds:
$v_{\theta}(x_{t},t)\approx\frac{-t\nabla_{x}\log p(x_{t})-x_{t}}{1-t}\;.$ (23)
Hence, we can approximate the score of the noisy prior with our learned velocity field $v_{\theta}$
$\nabla_{x_{t}}\log p(x_{t})\approx-\frac{(1-t)v_{\theta}(x_{t},t)+x_{t}}{t}\;,$ (24)
and we see that for $\omega(t)=\frac{t}{1-t}$ we obtain the conditional flow matching objective for $\mathcal{R}(x)$. We therefore set $\omega(t)=\frac{t}{1-t}$ and end up at our final objective:
$\arg\min_{q(x_{0}|y)}\underbrace{\mathbb{E}_{q(x_{0}|y)}\left[\frac{\|y-f(\mu_{x})\|^{2}}{2\nu^{2}}\right]}_{\text{$\mathcal{D}(x,y)$}}+\underbrace{\int_{0}^{T}~\mathbb{E}_{q(x_{t}|y)}\left[\left\|v_{\theta}(x_{t},t)-u_{t}(x_{t}|\epsilon)\right\|^{2}\right]dt}_{\text{$\mathcal{R}(x)$}}\;.$ (25)
Again, the gradient step for the regularizer becomes:
$\nabla_{\mu_{x}}\mathcal{R}(x)=\mathbb{E}_{t,q(x_{t}|y)}\left[v_{\theta}(x_{t},t)-u_{t}(x_{t}|\epsilon)\right]\;.$ (26)
A.2 Derivation of trajectory adjusted flow-based variational formulation
To achieve the proposed trajectory adjustment, we modify the forward process to:
$\hat{x}_{1}=x_{t+dt}+(1-t-dt)v_{\theta}(x_{t+dt},t+dt)\;,$ (27)
$x_{t}=(1-t)\mu_{x}+t\underbrace{(\alpha\hat{x}_{1}+\sqrt{1-\alpha^{2}}\epsilon)}_{\text{$\hat{\epsilon}$}}\;$, (28)
where $\hat{x}_{1}$ is the noise vector prediction from the last optimization iteration. This induces a variational distribution:
$q(x_{t}\mid y)=\mathcal{N}\left((1-t)\mu_{x}+t\alpha\hat{x}_{1},\,t^{2}(1-\alpha^{2})I\right)\;,$ (29)
leading to a score of
$\nabla_{x_{t}}\log q(x_{t}\mid y)=-\frac{1}{t^{2}(1-\alpha^{2})}\cdot t\sqrt{1-\alpha^{2}}\epsilon=-\frac{\epsilon}{t\sqrt{1-\alpha^{2}}}\;.$ (30)
The velocity field is again computed by Equation 18. We start by defining the flow:
$\psi_{t}(x_{0}\mid\epsilon)=(1-t)x_{0}+t\left(\alpha\hat{x}_{1}+\sqrt{1-\alpha^{2}}\epsilon\right)\;.$ (31)
The resulting derivative reads
$\frac{d}{dt}\psi_{t}(x_{0}\mid\epsilon)=\alpha\hat{x}_{1}-x_{0}+\sqrt{1-\alpha^{2}}\epsilon\;,$ (32)
and the inverse becomes
$x_{0}=\psi_{t}^{-1}(x_{t}\mid\epsilon)=\frac{x_{t}-t\alpha\hat{x}_{1}-t\sqrt{1-\alpha^{2}}\epsilon}{1-t}\;.$ (33)
Plugging these results into Equation 18:
$u_{t}(x_{t}\mid\epsilon)=\frac{\alpha\hat{x}_{1}+\sqrt{1-\alpha^{2}}\epsilon-x_{t}}{1-t}\;.$ (34)
A.3 Derivation of Score from Flow
The score matching objective reads as:
$\nabla_{x_{t}}\ln p_{t}(x_{t})=\underset{\theta}{\arg\min}~\mathbb{E}_{t\sim\mathcal{U}[0,1],x_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,I)}\left[w(t)\cdot\left\|s_{\theta}(x_{t},t)+\frac{1}{\sigma(t)^{2}}\left(x_{t}-\mu(x_{0},t)\right)\right\|^{2}\right],$ (35)
where,
$\frac{1}{\sigma(t)^{2}}\left(x_{t}-\mu(x_{0},t)\right)=\nabla_{x_{t}}\log p_{t}(x_{t}\mid x_{0}),$ (36)
with $p_{t}(x_{t}\mid x_{0})=\mathcal{N}(\mu_{t}(x_{0}),\sigma_{t}^{2}I)$. Note that as usual we assume $\mu_{t}(x_{0})$ being linear in $x_{0}$.
Equation 35 is solved by:
$\nabla_{x_{t}}\log p_{t}(x_{t})=\mathbb{E}_{p_{t}(x_{0}|x_{t})}\left[\nabla_{x_{t}}\log p_{t}(x_{t}|x_{0})\right],$ (37)
and can be written as:
$\nabla_{x_{t}}\log p_{t}(x_{t})=\frac{-(x_{t}-\mu(\mathbb{E}[x_{0}\mid x_{t}],t))}{\sigma(t)^{2}}.$ (38)
In the case of OT flow-matching, we obtain
$x_{t}=(1-t)x_{0}+tx_{1},$ (39)
$x_{1}\sim\mathcal{N}(0,\mathrm{Id})$ and $p(x_{t}|x_{0})=\mathcal{N}((1-t)x_{0},t^{2})$. The optimal velocity under the flow matching loss is given by:
$v^{*}(x_{t},t)=\mathbb{E}[x_{1}-x_{0}\mid x_{t}].$ (40)
Expressing $x_{1}=\frac{x_{t}-(1-t)x_{0}}{t}$, we can insert into Equation 40 and obtain:
$\mathbb{E}[x_{0}\mid x_{t}]=x_{t}-t\mathbb{E}[x_{1}-x_{0}\mid x_{t}].$ (41)
Inserting in Equation 38 leads to:
$\nabla_{x_{t}}\log p_{t}(x_{t})=-\frac{x_{t}-(1-t)(x_{t}-t\mathbb{E}[x_{1}-x_{0}\mid x_{t}])}{t^{2}},$ (42)
which for $v^{*}(x_{t},t)=\mathbb{E}[x_{1}-x_{0}\mid x_{t}]$ reads as:
$\nabla_{x_{t}}\log p(x_{t})\approx-\frac{(1-t)v_{\theta}(x_{t},t)+x_{t}}{t}.$ (43)
A.4 Implementation details
Flow Model and Regularizer Settings. As flow matching model, we us Stable Diffusion 3.5-Medium, which has been released under the Stability Community License. The classifier-free guidance scale is set to 2 for all experiments. To minimize the regularization term, we use stochastic gradient descent with a learning rate of $1$.
Data Likelihood Term. We use stochastic gradient descent for the minimization of the data term towards hard data consistency. For numerical stability, the squared error is summed over all measurements instead of computing the mean. The learning rate has to be adjusted accordingly, to compensate for the varying number of measurements $y$. Moreover, the minimization is terminated with early stopping once the likelihood term reaches $1\times 10^{-4}\cdot\text{len}(y)$, to not overfit the noise in the image observation.
Super-resolution. We employ bicubic downsampling as the forward operator, as implemented in [52]. The learning rate is set to 12 for $\times$12 super-resolution and to 6 for $\times$8 super-resolution.
Motion Deblurring. A different motion blur kernel is created for each sample using the MotionBlur package [8], available via github, with kernel size 61 and intensity 0.5. The learning rate for our data term optimizer is set to $10^{-1}$.
Inpainting. For inpainting on FFHQ we always use the same rectangular mask at a fixed position, chosen such that it roughly masks out the right side of the face (Figure 3). For DIV2k we also use a fixed mask for all samples, consisting of six randomly generated rectangles (Figure 6).
Data. We use the publicly available Flickr Faces High Quality dataset [23], which is realeased under the Creative Commons BY 2.0 License and the DIV2K dataset [3], which is released under a research only license. For FFHQ we use the first 1000 samples of the evaluation dataset and for DIV2K we use the 800 training samples. We downscale both datasets to $768\times 768$ px by applying bicubic sampling so that the shorter edge of the frame has 768 px and apply central cropping afterwards.
A.5 Baselines
For comparability, all baselines use Stable Diffusion 3.5-Medium and the same task definitions as in A.4.
FlowDPS [25] The standard FlowDPS implementation [26] is applied with 50 NFE, a classifier-free guidance scale of 2, and step sizes of 15 for inpainting and 10 for all other tasks.
FlowChef [38] Additionally, [26] is employed for FlowChef as well, using 200 NFE for inpainting and 50 NFE for all other tasks, a classifier-free guidance scale of 2, and a step size of 1 for all tasks.
Repulsive Score Distillation (RSD) [61]. We implement RSD for flow-matching models by applying Proposition 1 with $\omega(t)=t$, resulting in a weighting term consistent with the original RSD approach. However, we omit the pixel-space augmentation as it negatively affected performance when combined with the SD3 VAE. Consistent with the original findings from RSD, we observed that incorporating the repulsive term improves sample diversity but reduces fidelity. Therefore, we set the repulsive term to 0 for all results presented in the table, employing it exclusively for comparing posterior variances.
ReSample [47] We re implement ReSample for flow-matching by setting $\bar{\alpha}_{t}=\frac{\left(1-t\right)^{2}}{t^{2}+\left(1-t\right)^{2}}$. Furthermore, we compute the hard-data consistency at every iteration as larger skip steps seem to harm performance. We set the learning rate of the data term optimizer to 15 for all inverse problems.
PSLD [41] Our attempt to adapt PSLD following [26]—using 500 NFE, a classifier-free guidance scale of 2, and step sizes of 1 ($\times$12 super-resolution), 0.5 ($\times$8 super-resolution and motion deblurring) and 0.1 (inpainting)—did not yield meaningful results.
A.6 Regularizer weighting
Figure 1 displays the mean and standard deviation of the conditional flow matching loss $\mathcal{L}_{CFM}$ as a function of $t$, estimated over 100 samples. The loss function starts with high values at $t=1$, decreases over time, but then starts to rise again, and when reaching $t\approx 0.2$ even exceeds its initial value . The rising loss when approaching $t=0$ is due, in part, to the increasing difficulty of distinguishing high-frequency image content from residual noise. Another factor is that near $t=0$ the model operates in a highly sensitive regime where small prediction errors can cause disproportionately large deviations from the target, making accurate flow estimation particularly challenging in the final stages of the trajectory. We therefore modulate the regularization term according to the model error. Different weighting functions for $f(\mathcal{L}_{CFM})$ could be chosen that fulfill the condition $\lambda_{\mathcal{R}(t=0)}=0$. We simply take the reciprocal of the model error $\lambda_{\mathcal{R}(t)}=\mathcal{L}_{CFM,t}^{-1}$ as the regularization weight while $t\geq 0.2$, then set it to 0 for $t<0.2$. An alternative would be to shift the reciprocal of $\mathcal{L}_{CFM,t}^{-1}$ by $\mathcal{L}_{CFM,t=0}^{-1}$, such that $\lambda_{\mathcal{R}(t)}=\mathcal{L}_{CFM,t}^{-1}-\mathcal{L}_{CFM,t=0}^{-1}$. In Table 1 we compare our default weighting with this variant, denoted as $\lambda_{shift}$.
Figure 1: The Flow-Matching loss over time $t$.
Table 1: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$. We compare different weighting functions $\lambda_{\mathcal{R}(t)}$ based on the model error
SR $\times$8 SR $\times$12 Motion Deblurring Inpainting
[…]
FFHQ $\times768$
$\lambda_{shift}$ $0.246$ $27.4$ $0.793$ $29.91$ $0.286$ $24.4$ $0.766$ $28.19$ $0.237$ $14.5$ $0.790$ $29.84$ $0.180$ $8.2$ $0.828$ $23.58$
Ours 0.213 235 303 759 574 90.213\,235\,303\,759\,574\,9 13.266 211 509 704 5913.266\,211\,509\,704\,59 $0.777$ $29.54$ 0.270 564 854 145 050 050.270\,564\,854\,145\,050\,05 16.201 229 095 458 98416.201\,229\,095\,458\,984 0.740 209 460 258 483 90.740\,209\,460\,258\,483\,9 27.714 468 002 319 33627.714\,468\,002\,319\,336 0.236 155 018 210 411 070.236\,155\,018\,210\,411\,07 10.732 836 723 327 63710.732\,836\,723\,327\,637 $0.772$ 29.611 394 882 202 1529.611\,394\,882\,202\,15 0.183 834 627 270 698 550.183\,834\,627\,270\,698\,55 8.741 712 570 190 438.741\,712\,570\,190\,43 0.827 794 373 035 430 90.827\,794\,373\,035\,430\,9 23.693 054 199 218 7523.693\,054\,199\,218\,75
DIV2K $\times768$
$\lambda_{shift}$ $0.379$ $30.4$ $0.625$ $23.58$ $0.434$ $37.5$ $0.522$ $21.40$ $0.337$ $25.8$ $0.664$ $24.54$ $0.151$ $9.0$ $0.819$ $23.79$
Ours 0.352 998 495 101 928 70.352\,998\,495\,101\,928\,7 26.456 972 122 192 38326.456\,972\,122\,192\,383 $0.607$ $23.30$ 0.421 490 550 041 198 730.421\,490\,550\,041\,198\,73 32.118 663 787 841 832.118\,663\,787\,841\,8 0.525 365 948 677 0630.525\,365\,948\,677\,063 21.385 475 158 691 40621.385\,475\,158\,691\,406 0.314 553 380 012 512 20.314\,553\,380\,012\,512\,2 21.111 848 831 176 75821.111\,848\,831\,176\,758 0.652 780 890 464 782 70.652\,780\,890\,464\,782\,7 24.444 496 154 785 15624.444\,496\,154\,785\,156 0.162 881 195 545 196 530.162\,881\,195\,545\,196\,53 11.026 673 316 955 56611.026\,673\,316\,955\,566 0.814 652 204 513 549 80.814\,652\,204\,513\,549\,8 23.752 101 898 193 3623.752\,101\,898\,193\,36
A.7 Effect of captioning
Given the diversity of DIV2k, we use DAPE [56] to generate captions for it and include them in the prompt A high quality photo of [DAPE caption]. For FFHQ we always prompt with A high quality photo of a face.. The effect of the text prompt is to increase the likelihood of our sample under the prior of the (pre-trained, frozen) image generator. For comparison, we also ran experiments without data specific captions, where we always used the generic prompt A high quality photo. Results are shown in Table 2
Table 2: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$. We compare our version with data-specific captions and a version without captions.
FFHQ DIV2K
[…]
wo captions 0.278 17.0 0.734 27.66 0.488 51.5 0.546 21.82
Ours 0.271 16.2 0.740 27.71 0.421 32.1 0.525 21.39
A.8 Additional Experimental Results
Table 3: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$ – Super-resolution ($\times 8$ and $\times 12$).
SR $\times 8$ SR $\times 12$
[…]
FFHQ $\times768$
ReSample 0.400 $\pm$ 0.069 55.6 0.815 $\pm$ 0.051 26.37 $\pm$ 1.00 0.474 $\pm$ 0.078 80.3 0.786 $\pm$ 0.056 25.47 $\pm$ 1.16
FlowDPS 0.374 $\pm$ 0.107 38.5 0.756 $\pm$ 0.075 29.24 $\pm$ 2.04 0.413 $\pm$ 0.107 44.0 0.741 $\pm$ 0.074 28.05 $\pm$ 2.06
RSD 0.391 $\pm$ 0.079 51.7 0.776 $\pm$ 0.052 29.69 $\pm$ 2.04 0.462 $\pm$ 0.093 71.7 0.743 $\pm$ 0.059 28.11 $\pm$ 2.00
FlowChef 0.341 $\pm$ 0.083 30.5 0.760 $\pm$ 0.064 28.42 $\pm$ 2.22 0.373 $\pm$ 0.084 46.5 0.730 $\pm$ 0.068 27.00 $\pm$ 2.07
Ours 0.213 $\pm$ 0.056 13.3 0.777 $\pm$ 0.051 29.54 $\pm$ 2.02 0.271 $\pm$ 0.071 16.2 0.740 $\pm$ 0.058 27.71 $\pm$ 2.00
DIV2K $\times768$
ReSample 0.533 $\pm$ 0.130 55.0 0.625 $\pm$ 0.132 22.34 $\pm$ 2.27 0.643 $\pm$ 0.152 88.1 0.562 $\pm$ 0.151 20.85 $\pm$ 3.02
FlowDPS 0.476 $\pm$ 0.129 44.4 0.567 $\pm$ 0.139 23.01 $\pm$ 3.01 0.547 $\pm$ 0.139 54.0 0.528 $\pm$ 0.146 21.79 $\pm$ 2.94
RSD 0.539 $\pm$ 0.121 60.9 0.591 $\pm$ 0.124 23.45 $\pm$ 2.96 0.684 $\pm$ 0.137 95.7 0.523 $\pm$ 0.132 21.96 $\pm$ 2.86
FlowChef 0.490 $\pm$ 0.116 36.5 0.539 $\pm$ 0.137 21.84 $\pm$ 2.96 0.525 $\pm$ 0.118 43.8 0.492 $\pm$ 0.145 20.52 $\pm$ 2.85
Ours 0.353 $\pm$ 0.112 26.5 0.607 $\pm$ 0.127 23.30 $\pm$ 2.90 0.421 $\pm$ 0.131 32.1 0.525 $\pm$ 0.136 21.39 $\pm$ 2.67
Table 4: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$ – Motion deblurring and in-painting.
Motion Deblurring In-painting
[…]
FFHQ $\times768$
ReSample 0.457 $\pm$ 0.087 82.9 0.788 $\pm$ 0.058 25.45 $\pm$ 1.46 0.366 $\pm$ 0.053 70.8 0.827 $\pm$ 0.033 21.83 $\pm$ 1.68
FlowDPS 0.431 $\pm$ 0.117 54.3 0.732 $\pm$ 0.078 27.64 $\pm$ 2.20 0.344 $\pm$ 0.060 42.5 0.771 $\pm$ 0.048 19.19 $\pm$ 3.19
RSD 0.458 $\pm$ 0.098 77.3 0.743 $\pm$ 0.059 27.67 $\pm$ 2.47 0.478 $\pm$ 0.082 73.3 0.736 $\pm$ 0.048 21.97 $\pm$ 2.58
FlowChef 0.406 $\pm$ 0.093 40.2 0.716 $\pm$ 0.072 25.81 $\pm$ 2.61 0.394 $\pm$ 0.069 69.8 0.780 $\pm$ 0.051 18.18 $\pm$ 2.84
Ours 0.236 $\pm$ 0.070 10.7 0.772 $\pm$ 0.055 29.61 $\pm$ 2.24 0.184 $\pm$ 0.038 8.7 0.828 $\pm$ 0.029 23.69 $\pm$ 2.77
DIV2K $\times768$
ReSample 0.556 $\pm$ 0.146 79.7 0.617 $\pm$ 0.134 21.79 $\pm$ 2.52 0.285 $\pm$ 0.073 51.9 0.796 $\pm$ 0.067 22.68 $\pm$ 1.84
FlowDPS 0.558 $\pm$ 0.153 65.5 0.536 $\pm$ 0.148 21.88 $\pm$ 3.02 0.328 $\pm$ 0.103 29.2 0.692 $\pm$ 0.112 21.71 $\pm$ 2.67
RSD 0.638 $\pm$ 0.156 97.6 0.551 $\pm$ 0.136 22.10 $\pm$ 3.07 0.464 $\pm$ 0.112 63.9 0.678 $\pm$ 0.077 23.23 $\pm$ 2.21
FlowChef 0.561 $\pm$ 0.123 49.6 0.486 $\pm$ 0.148 19.90 $\pm$ 3.06 0.489 $\pm$ 0.148 58.3 0.659 $\pm$ 0.131 20.87 $\pm$ 2.65
Ours 0.315 $\pm$ 0.107 21.1 0.653 $\pm$ 0.121 24.44 $\pm$ 3.05 0.163 $\pm$ 0.053 11.0 0.815 $\pm$ 0.054 23.75 $\pm$ 2.74
We present the experimental results from the main paper in Table 3, now augmented with sample-wise standard deviations for all metrics except FID.
A.9 FLAIR in Pixel Space
Table 5: Quantitative results with 50 NFE and $\sigma_{\nu}=0.5\%$ – In-painting and Super-resolution ($\times 8$).
Inpainting SR $\times 8$
[…]
DDNM 0.158 $\pm$ 0.042 26.9 0.732 $\pm$ 0.037 18.31 $\pm$ 2.94 0.199 $\pm$ 0.052 31.9 0.635 $\pm$ 0.079 23.59 $\pm$ 1.64
DPS 0.195 $\pm$ 0.064 30.2 0.689 $\pm$ 0.077 20.49 $\pm$ 2.81 0.172 $\pm$ 0.058 27.8 0.658 $\pm$ 0.088 24.59 $\pm$ 2.04
$M$ 0.161 $\pm$ 0.054 28.8 0.728 $\pm$ 0.062 20.59 $\pm$ 3.37 0.172 $\pm$ 0.051 29.1 0.669 $\pm$ 0.083 24.65 $\pm$ 1.97
[…]
FLAIR 0.097 $\pm$ 0.035 14.2 0.831 $\pm$ 0.031 21.87 $\pm$ 2.66 0.143 $\pm$ 0.039 22.9 0.712 $\pm$ 0.076 25.93 $\pm$ 1.96
[…]
Figure 2: Qualitative comparison. FLAIR in pixel space produces posterior samples of high perceptual quality while maintaining high data likelihood as well. Best viewed zoomed in.
[…]
A.10 Runtime Analysis
We compare the runtime and memory consumption of our method to the baselines. As our hard data consistency can strongly influence the runtime, we also provide measurements with the number of data term steps $\leq 5$ and additionally a fast version using a "tinyVAE" of SD3. To validate that the usage of the tinyVAE or less steps does not degrades the performance noticeably we also provide a metrics for x12 Super Resolution on 100 samples of FFHQ:
Table 6: Comparison of different methods in terms of runtime and memory usage. We validate the use of less data steps and a "tinyVAE" on $\times 12$ super resolution on 100 samples of the FFHQ dataset.
Method Runtime (s) $\downarrow$ Memory (MB) $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$
Resample 88.02 19009.2 0.461 25.31
FlowDPS 34.15 12228.6 0.404 27.74
RSD (no repulsion) 21.19 12400.0 0.462 28.11
FlowChef 15.23 12227.72 0.361 26.56
FLAIR (HDC, large VAE) 172.34 12389.4 0.259 27.42
FLAIR (HDC, tiny VAE) 40.77 5960.2 0.256 27.59
FLAIR (5 data term steps, tiny VAE) 22.46 5960.2 0.264 27.61
A.11 Ensembling Experiment
[…]
For ensembling we averaged 8 independent predictions of the corresponding methods. It can be seen that PSNR improves for all methods when ensembling. However, FLAIR shows the biggest gain, which means that our samples are indeed distributed around the mean and feature a higher variance compared to the baselines.
SR $\times 8$ SR $\times 12$
[…]
DIV2K
FlowDPS 22.53 0.4837 21.47 0.5524
FlowDPS (8x ensemble) 23.28 0.5157 22.06 0.5995
FlowChef 21.44 0.4898 20.20 0.5228
FlowChef (8x ensemble) 22.60 0.5502 21.60 0.5931
FLAIR 22.83 0.3627 21.05 0.4270
FLAIR (8x ensemble) 23.79 0.4244 22.27 0.4930
FFHQ
FlowDPS 29.02 0.3659 27.74 0.4036
FlowDPS (8x ensemble) 30.12 0.3267 28.65 0.3749
FlowChef 28.05 0.3303 26.56 0.3609
FlowChef (8x ensemble) 29.54 0.3267 28.15 0.3602
FLAIR 29.36 0.2028 27.42 0.2594
FLAIR (8x ensemble) 30.94 0.2457 29.00 0.2999
To highlight that our model can also be used to obtain good MMSE estimates, we also conducted ensemble predictions by running posterior sampling 8 times and averaging the result. The results show that ensembling increases PSNR values, but reduces LPIPS and confirms that our samples are indeed distributed around the posterior mean and that results very close to the posterior mean like the baseline methods are perceptually further away (LPIPS) from the ground truth compared to our samples.
A.12 Statistical Relevance
Our method is training-free, and the variance in reconstructed images is intentional, reflecting the stochasticity of our sampling process rather than instability. All methods are evaluated with identical random seeds to ensure fair comparison. We compute metrics over 1000 samples for FFHQ and 800 samples for DIV2K. Perceptual FID (pFID) is evaluated on $256\times 256$ patches, resulting in 9000 and 7200 samples, respectively.  Table 3 in Appendix A.8 reports means and standard deviations over multiple samples.
We also evaluated 100 FFHQ samples and 80 DIV2K samples, sampling three reconstructions per input for each method. We report the mean of each metric across all samples and the standard deviation of the means.
Table 8: Statistical evaluation on FFHQ for $\times$8 Super Resolution. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.370 $\pm$ 0.0012 70.7 $\pm$ 1.45 0.755 $\pm$ 0.0010 28.98 $\pm$ 0.008
RSD 0.4678 $\pm$ 0.0001 102.9 $\pm$ 0.05 0.7362 $\pm$ 0.0001 28.45 $\pm$ 0.001
FlowChef 0.3316 $\pm$ 0.0055 63.5 $\pm$ 1.45 0.7593 $\pm$ 0.0027 28.12 $\pm$ 0.072
Ours 0.2039 $\pm$ 0.0048 40.5 $\pm$ 0.94 0.7970 $\pm$ 0.0228 29.74 $\pm$ 0.668
Table 9: Statistical evaluation on FFHQ for $\times$12 Super Resolution. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.4073 $\pm$ 0.0002 77.6 $\pm$ 1.05 0.7391 $\pm$ 0.0006 27.71 $\pm$ 0.016
RSD 0.5039 $\pm$ 0.0002 119.0 $\pm$ 0.12 0.7217 $\pm$ 0.0001 27.08 $\pm$ 0.001
FlowChef 0.3626 $\pm$ 0.0050 81.1 $\pm$ 1.03 0.7283 $\pm$ 0.0027 26.62 $\pm$ 0.059
Ours 0.2593 $\pm$ 0.0023 45.8 $\pm$ 1.51 0.7582 $\pm$ 0.0252 27.81 $\pm$ 0.660
Table 10: Statistical evaluation on FFHQ for Motion Blur. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.4140 $\pm$ 0.0030 83.83 $\pm$ 1.00 0.7383 $\pm$ 0.0006 27.47 $\pm$ 0.05
RSD 0.4515 $\pm$ 0.0001 108.75 $\pm$ 0.08 0.7437 $\pm$ 0.0001 27.40 $\pm$ 0.00
FlowChef 0.4019 $\pm$ 0.0007 74.89 $\pm$ 0.69 0.7178 $\pm$ 0.0022 25.50 $\pm$ 0.04
Ours 0.2196 $\pm$ 0.0080 38.8 $\pm$ 2.25 0.7964 $\pm$ 0.0319 30.10 $\pm$ 0.96
Table 11: Statistical evaluation on FFHQ for Inpainting. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.3315 $\pm$ 0.0015 74.00 $\pm$ 0.58 0.7755 $\pm$ 0.0007 19.06 $\pm$ 0.11
RSD 0.4601 $\pm$ 0.0003 103.02 $\pm$ 0.03 0.7430 $\pm$ 0.0000 22.19 $\pm$ 0.01
FlowChef 0.3771 $\pm$ 0.0013 102.22 $\pm$ 0.26 0.7888 $\pm$ 0.0016 18.42 $\pm$ 0.25
Ours 0.1761 $\pm$ 0.0012 33.23 $\pm$ 2.02 0.8423 $\pm$ 0.0172 24.07 $\pm$ 0.80
Table 12: Statistical evaluation on DIV2K for $\times 8$ Super Resolution. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.5517 $\pm$ 0.0046 138.24 $\pm$ 1.67 0.5207 $\pm$ 0.0022 22.41 $\pm$ 0.02
RSD 0.7163 $\pm$ 0.0002 181.83 $\pm$ 0.15 0.4892 $\pm$ 0.0001 21.99 $\pm$ 0.00
FlowChef 0.5726 $\pm$ 0.0046 145.77 $\pm$ 3.31 0.4998 $\pm$ 0.0014 21.21 $\pm$ 0.04
Ours 0.3716 $\pm$ 0.0161 88.08 $\pm$ 1.43 0.5991 $\pm$ 0.0192 23.06 $\pm$ 0.41
Table 13: Statistical evaluation on DIV2K for $\times 12$ Super Resolution. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.6264 $\pm$ 0.0045 154.31 $\pm$ 0.43 0.4866 $\pm$ 0.0024 21.37 $\pm$ 0.02
RSD 0.7714 $\pm$ 0.0002 198.85 $\pm$ 0.12 0.4683 $\pm$ 0.0001 21.15 $\pm$ 0.00
FlowChef 0.6020 $\pm$ 0.0060 151.32 $\pm$ 2.45 0.4586 $\pm$ 0.0020 20.10 $\pm$ 0.03
Ours 0.4316 $\pm$ 0.0151 101.12 $\pm$ 4.51 0.5236 $\pm$ 0.0229 21.35 $\pm$ 0.51
Table 14: Statistical evaluation on DIV2K for Motion Deblur. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.6242 $\pm$ 0.0082 161.57 $\pm$ 3.90 0.4978 $\pm$ 0.0028 21.41 $\pm$ 0.04
RSD 0.8067 $\pm$ 0.0002 216.37 $\pm$ 0.36 0.4364 $\pm$ 0.0001 20.82 $\pm$ 0.00
FlowChef 0.6292 $\pm$ 0.0007 158.08 $\pm$ 3.03 0.4557 $\pm$ 0.0025 19.62 $\pm$ 0.05
Ours 0.3069 $\pm$ 0.0036 77.77 $\pm$ 1.89 0.6596 $\pm$ 0.0316 24.46 $\pm$ 0.71
Table 15: Statistical evaluation on DIV2K for Inpainting. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.3738 $\pm$ 0.0032 106.58 $\pm$ 0.89 0.6579 $\pm$ 0.0009 21.06 $\pm$ 0.05
RSD 0.4667 $\pm$ 0.0003 136.82 $\pm$ 0.21 0.6650 $\pm$ 0.0001 23.07 $\pm$ 0.00
FlowChef 0.5111 $\pm$ 0.0019 128.97 $\pm$ 0.60 0.6355 $\pm$ 0.0006 20.51 $\pm$ 0.02
Ours 0.1729 $\pm$ 0.0014 51.41 $\pm$ 0.48 0.8122 $\pm$ 0.0124 24.06 $\pm$ 0.87
A.12.1 t-Test Analysis
We further performed paired t-tests on the LPIPS scores between FlowDPS and FLAIR. The null hypothesis states that the mean LPIPS scores are the same for both methods. In all settings, we reject the null hypothesis ($p<0.001$), confirming the statistical significance of our improvements see Table 16.
Table 16: Paired t-test $p$-values for LPIPS (FlowDPS vs. Ours). All comparisons are statistically significant.
Dataset Task p-value
[…]
Motion Deblur $5.30\times 10^{-4}$
Inpainting $1.21\times 10^{-4}$
[…]
Motion Deblur $4.29\times 10^{-5}$
Inpainting $3.00\times 10^{-5}$
A.13 Additional Qualitative Examples
To illustrate the visual differences behind the error metrics, we present additional qualitative results for both FFHQ and DIV2k, comparing FLAIR with existing approaches. These examples complement the images in the main paper and highlight the visual fidelity, consistency, and robustness of our method across diverse scenes and different degradations. Figure 9 features a full sized version of the variance figure in section subsection 5.2.
Figure 3: Inpainting results on FFHQ. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces realistic, high-frequency details while previous works either fail to inpaint the region correctly or collapse to overly smooth solutions.
Figure 4: $\times 12$ super-resolution results on FFHQ. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
Figure 5: Motion de-blur results on FFHQ. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
Figure 6: Inpainting results on DIV2k. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces realistic, high-frequency details while previous works either fail to inpaint the region correctly or collapse to overly smooth solutions. Moreover they do not fit the data term (not inpainted region) very well.
Figure 7: $\times 12$ super-resolution results on DIV2k. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
Figure 8: Motion de-blur results results on DIV2k. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
Refer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSD
Figure 9: Individual samples for x12 Super Resolution with zoom and std. FLAIR produces varied samples from the posterior. For superresoltion The variance is expected to be mostly in the high frequencies, because the data term limits low frequency variations. The baselines tend to predict very similar looking images with less detail.
A.14 Failure cases
We observe two main failure modes for FLAIR, see Figure 10. First, we find that super-resolution on DIV2k occasionally results in grainy textures, usually in regions with abundant high-frequency detail and complicated light transport. Potentially, this happens for images which do not have high probability under the prior. We do not observe those artifacts for the FFHQ dataset. Second, we observe a few instances where the strong generative prior hallucinates semantically inconsistent or misaligned structures – especially facial features.
Figure 10: Qualitative failure cases of FLAIR on DIV2k and FFHQ. Top row: grainy results from systematic error. Those errors potentially stem from a weak prior for those images. For example we do not observe them for the FFHQ dataset Bottom row: Semantically inconsistent failures. Sometimes the model lacks the ability to incorporate globally consistent semantics into its restorations.